### v1

# Execution-Feedback Driven Test Generation from SWE Issues

Abstract.
A software engineering issue (SWE issue) is easier to resolve when accompanied by a reproduction test. Unfortunately, most issues do not come with functioning reproduction tests, so this paper explores how to generate them automatically. The primary challenge in this setting is that the code to be tested is either missing or wrong, as evidenced by the existence of the issue in the first place. This has held back test generation for this setting: without the correct code to execute, it is difficult to leverage execution feedback to generate good tests. This paper introduces novel techniques for leveraging execution feedback to get around this problem, implemented in a new reproduction test generator called e-Otter++. Experiments show that e-Otter++ represents a leap ahead in the state-of-the-art for this problem, generating tests with an average fail-to-pass rate of 63% on the TDD-Bench Verified benchmark.
LLMs, SWE Patches, Reproduction Tests
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Software engineering (SWE) issues are descriptions of a bug or missing feature. SWE issues are usually informal and lack automated tests. But such tests, when they exist, are invaluable for software development. Before issue resolution, tests can reproduce the issue to confirm whether it is real, make it more concrete, and help stakeholders agree upon the desired behavior—following the highly regarded software development practice of test-driven development (TDD) (Beck, 2002). Once a code change (a patch) is proposed to resolve the issue, running the same reproduction tests helps decide whether to accept the patch. Finally, after issue resolution, adding the tests to the project’s test suite helps prevent future regressions. Reproduction tests for SWE issues should be fail-to-pass ($F\!\!\to\!\!P$): they should fail on the old (pre-patch) code to reproduce the issue and pass on the new (post-patch) code to accept the patch. Unfortunately, generating $F\!\!\to\!\!P$ tests is challenging, because issue descriptions are informal and because the new code patch on which the tests should pass does not yet exist. In addition to helping human software engineers, generating tests from SWE issues can also assist automated SWE agents, as several SWE agents also rely on reproduction tests (Arora et al., 2024; Chen et al., 2024; Ehrlich et al., 2025; Jain et al., 2025; Li et al., 2025; Ruan et al., 2024; Xia et al., 2025; Yang et al., 2024; Zhang et al., 2024).
There have been several recent attempts at using large language models (LLMs) to automatically generate $F\!\!\to\!\!P$ tests from SWE issues, without access to the resolving code patches. Libro uses an LLM to directly generate 50 tests from Java bug reports, then performs rule-based postprocessing and ranking (Kang et al., 2023). Other attempts focus on Python, modifying a SWE agent designed to automatically generate code patches to generate tests instead (Lin et al., 2024; Mündler et al., 2024). Recent papers proposed novel LLM-based flows or agents custom-tailored to the problem of generating tests from issues, including AEGIS (Wang et al., 2024), Otter++ (Ahmed et al., 2025), BRT agent (Cheng et al., 2025), and Issue2Test (Nashid et al., 2025).
There are two benchmarks for generating tests from issues: SWT-bench (Mündler et al., 2024) (derived from SWE-bench (Jimenez et al., 2024)) and TDD-Bench Verified (Ahmed et al., 2024) (derived from SWE-bench Verified (Chowdhury et al., 2024)). SWT-bench Verified, also derived from SWE-bench Verified, is very similar to TDD-Bench Verified. SWT-bench has a leaderboard (https://swtbench.com), topped by the unpublished Amazon Q Developer Agent with an $F\!\!\to\!\!P$ rate of 37.7% on SWT-bench Lite and 49.0% on SWT-bench Verified. However, as there is no paper about Amazon Q Developer Agent, its inner workings are not publicly known. The top system on SWT-bench Verified with a published paper is Otter++ (Ahmed et al., 2025) with an $F\!\!\to\!\!P$ rate of 37.0%, and it also has the exact same $F\!\!\to\!\!P$ rate of 37.0% on TDD-Bench Verified. Given how similar the two verified benchmarks are, this paper evaluates on TDD-Bench Verified only. This paper establishes a new state of the art (SOTA) for SWT-bench Lite with 52.5% $F\!\!\to\!\!P$ rate and for TDD-Bench Verified with 63.0% $F\!\!\to\!\!P$ rate. It does so with a new system, e-Otter++. The “e-” refers to execution feedback: e-Otter++ outperforms prior work by leveraging execution feedback more effectively.
The first place where e-Otter++ uses execution feedback is after generating a candidate test. The test should fail on the old code, which is easy to accomplish, but it should fail for the right reason, which is challenging. Unlike prior work, e-Otter++ uses execution feedback to retrieve related code and augments the repair prompt with that. The second place where e-Otter++ uses execution feedback is for test selection. After generating and repairing multiple tests, it automatically picks a single final test. The chosen test should pass on the new code, but since the resolving new code is not available, there is no execution feedback to tell whether that is the case. Unlike prior work, e-Otter++ generates candidate code patches as a surrogate for the hidden ground-truth new code. Despite generated code patches being of course imperfect, our experiments show that tests passing on them helps predict passing on the ground-truth code patch. However, even the best test selector is only as good as the pool of generated candidate tests it picks from. Ideally, this test pool should be diverse, and most prior work turns up the LLM decoding temperatures to increase candidate diversity. Unlike prior work, e-Otter++ instead gets diversity from two kinds of systematic prompt variations: issue description morphs, which use an LLM to rewrite the issue text in various ways before test generation, and prompt context masks, which elide various subsets of features from the test-generation prompt. Our experiments show that morphs and masks greatly increase the likelihood of at least one candidate test being $F\!\!\to\!\!P$.
This paper makes the following novel contributions:
  • Generating reproduction tests from issues and increasing their correctness via execution-augmented test repair.
  • Increasing reproduction test candidate diversity via issue description morphs and test-generation context masks.
  • Selecting a single final reproduction test via execution feedback from imperfect code patches.
Taken together, these novel contributions enable e-Otter++ to substantially advance the SOTA for generating reproduction tests from issues. Ultimately, we hope our work helps cement the practice of using tests both to drive issue resolution and for long-term continuous quality assurance after.
2. Problem Statement and Background
This section discusses the problem statement and briefly introduces prior work we build upon: Otter (Ahmed et al., 2025), Agentless (Xia et al., 2025), SWT-bench (Mündler et al., 2024), and TDD-Bench Verified (Ahmed et al., 2024).
Figure 1. Evaluation harness for bug reproduction test. (1) The issue description and $c_{\textrm{old}}$ go into e-Otter++ as input, (2) e-Otter++ generates a test patch with a reproduction test $y$, (3) executing the test $y$ on $c_{\textrm{old}}$ should fail in order to reproduce the issue, (4) a developer-written golden code patch is applied to $c_{\textrm{old}}$ as a pull request resulting in $c_{\textrm{new}}$, (6) executing the test $y$ on $c_{\textrm{new}}$ should pass now in order to confirm that the issue has been addressed by the pull request.
2.1. Problem Statement
This paper addresses the challenge of generating reproduction tests directly from issue descriptions. Figure 1 presents an evaluation harness for reproduction tests. The input $x$ consists of two parts: an issue description $d_{\textrm{issue}}$ (usually written in natural language and possibly including code snippets and stack traces) and the existing version of the codebase $c_{\textrm{old}}$ (before any fix is applied). Crucially, a solution for this problem statement does not have access to the updated codebase $c_{\textrm{new}}$, reflecting real-world conditions. The goal is to produce a set of tests $y$ that will fail on $c_{\textrm{old}}$, thereby confirming the presence of the issue, and pass on $c_{\textrm{new}}$, confirming that the issue has been resolved (in other words, $y$ should be $F\!\!\to\!\!P$). These tests should aim to cover the relevant changes made in the code. The problem is to develop a function genTests that takes as input only $x=\langle d_{\textrm{issue}},c_{\textrm{old}}\rangle$ and returns the generated tests $y$ as output.
2.2. Reproduction Test Generation: Otter
This paper introduces e-Otter++, which extends an earlier system called Otter++ (Ahmed et al., 2025). Both e-Otter++ and Otter++ generate tests directly from issue descriptions and $c_{\textrm{old}}$. In other words, both are functions genTests that address the problem stated in Section 2.1.
At the heart of Otter++ is Otter, which is a pipeline of LLM-based and rule-based steps. Otter comprises three key components: a localizer, a self-reflective action planner, and a test generator. The localizer identifies relevant files, test functions, and focal functions by querying the LLM with context-aware prompts. The action planner then iteratively builds a plan of actions—categorized as read, write, and modify—guided by a self-reflective loop and validation checks. Finally, the test generator produces complete test functions using localized code and structural cues like test signatures and imports; it also includes mechanisms for repairing hallucinated or incomplete imports and integrates static analysis tools (e.g., the Flake8 Python linter) to ensure syntactic and semantic correctness.
Otter++ improves Otter by generating multiple test candidates using heterogeneous prompts, each varying in the inclusion of localized context. It selects the best candidate by analyzing execution outcomes (e.g., assertion failures vs. syntax errors) on the original code $c_{\textrm{old}}$. Empirical results using two models (GPT-4o and Mistral-large) show that Otter++ achieves higher $F\!\!\to\!\!P$ rates than prior systems, while remaining cost-efficient and generalizable.
Our new system e-Otter++ uses Otter as a base component and adds several new techniques, discussed in detail in Section 3.
2.3. Code Patch Generation: Agentless
One of the new techniques in e-Otter++ is how it uses execution feedback from automatically generated candidate code patches. Rather than implementing a new code patch generator, it reuses an existing one, Agentless (Xia et al., 2025), as a component.
Agentless implements a hierarchical bug localization process. First, it identifies suspicious files using both LLM prompting and embedding-based retrieval (filtered to avoid irrelevant files). It then narrows down to relevant classes/functions using a concise skeleton representation of files. Finally, it pinpoints exact edit locations with line-level granularity. For repair, it generates multiple patches in a lightweight Search/Replace diff format rather than rewriting entire code blocks. This minimizes hallucinations and boosts cost-efficiency. During patch validation, Agentless automatically generates reproduction tests (when none are provided) and combines them with regression tests to select the best patch via majority voting. This modular structure allows Agentless to stay interpretable, easily debuggable, and effective without relying on complex environment interaction or planning mechanisms typical of agents.
Our new system e-Otter++ uses only the first part of Agentless, without the test generation or patch selection of Agentless, as a component (see Section 3.2). While we could have adopted any code patch generator, we chose Agentless because it is open-source and candidate patches are publicly available [^1] https://github.com/OpenAutoCoder/Agentless/tree/main.
2.4. Test Evaluation Benchmarks:
TDD-Bench Verified and SWT-bench Lite
There are two benchmarks for evaluating bug reproduction tests. SWT-bench Lite (Mündler et al., 2024) was derived primarily from the SWE-bench Lite dataset (Jimenez et al., 2024) and contains 276 samples. TDD-Bench Verified (Ahmed et al., 2024) was derived from SWE-bench Verified (Chowdhury et al., 2024). In TDD-Bench Verified, the authors ran 500 samples from SWE-bench Verified, filtered out samples with zero test coverage or those violating the $F\!\!\to\!\!P$ property of the tests, and ended up with 449 samples. Both benchmarks judge whether a generated test reproduces the described issue by checking whether the test fails before applying the golden patch and passes afterward. In other words, they both implement an evaluation harness consistent with the diagram in Figure 1. One difference between their implementations is that TDD-Bench only runs contributing tests (tests that were added or updated), whereas SWT-bench runs the whole test files that contain the contributing tests. Later, Mündler et al. introduced SWT-bench Verified, which also derives from SWE-bench Verified like TDD-Bench, into their leaderboard (https://swtbench.com). Section 4 reports results on both SWT-bench Lite and TDD-Bench Verified.
3. Methodology
Figure 2 presents the overview of our approach. It has three major components: test generator, code patch generator, and test selector. The following three subsections describe each component.
Figure 2. Overview of our approach.
3.1. Test Generator
Our test generator, e-Otter, starts with tests generated by Otter (see Section 2.2), then increases their correctness via execution-augmented test repair. It also increases reproduction test candidate diversity via issue description morphs and test-generation context masks.
Execution-augmented Test Repair.
The test repair component starts by running an Otter-generated candidate test $y$ on $c_{old}$. It is expected that the test should fail in this phase. However, a test may not fail for the reason mentioned in the issue description. Chen et al. (Chen et al., 2025) reported that Python execution errors during the issue resolution phase correlate with lower resolution rates. They also identified the most prevalent errors—such as ModuleNotFoundError and TypeError—in the generated code, which tend to differ greatly from the reason mentioned in the issue description. To address this, we propose an LLM-based critic that will take the issue description and test log as input and decide whether the test is failing for the right reason. If the critic believes the test is failing for the reason mentioned in the issue description, the repair loop immediately stops. Otherwise, it tries to repair the error with additional information identified by the critic, which includes the possible buggy line, the most relevant text/code snippet from the issue description, and any function that the model wants to see before attempting to write the test again. The critic consists of a single LLM call with temperature 0, where the LLM output contains both the decision and additional information. We explicitly ask for the buggy line because these Python execution errors (e.g., ModuleNotFoundError, TypeError) are localized to one particular line. It is helpful for the model to know which part of the program needs more attention. We have also seen some cases where the issue description has enough hints to avoid execution errors. For example, in “astropy__astropy-14182”, the GPT-4o model generates “ascii.write()”, which is wrong. The correct version is “tbl.write()”, which is not as highly used as “ascii.write()” but is present in the issue description. The critic also proposes function names to read to augment the context (e.g., function arguments, return type), and that actually helps the model in repairing the tests.
After obtaining the additional information from the critic, the repair loop collects the functions from the code that the LLM wants to see. Next, it calls the LLM again to repair the test, feeding the issue description, the current version of the test, the test log, and the additional information collected based on the critic output into the test-repair prompt. Then, the repair loop starts over from this new test as $y$, until the critic is satisfied or for a maximum of 10 iterations. Following prior works (Xia et al., 2025; Ehrlich et al., 2025; Chen et al., 2021), the test repair LLM call uses a higher temperature (0.8) to encourage the model to try diverse solutions. We refer to the test generated at this phase as e-Otter test.
(a) Issue description
(b) Test patch
Figure 3. Issue description and test patch for sympy__sympy-23413. Some lines are omitted from the patch due to space constraints.
Issue Description Morphs and Test-Generation Context Masks
Recent works both on generating SWE-patches (Ehrlich et al., 2025; Xia et al., 2025; Jain et al., 2025) and reproduction tests (Ahmed et al., 2025; Wang et al., 2024) have adopted inference scaling. Most works rely on higher temperatures to increase the diversity of solutions. One of the key limitations of such approaches is using the same context for all solutions, which may contain misleading information (e.g., incorrect localization). Also, in a chain of multiple LLM calls, using a higher temperature in each call exponentially increases the number of samples.
Instead of higher temperature, Otter++ (Ahmed et al., 2025) attempted heterogeneous prompting via five masks hiding different parts of contexts and achieved a higher fail-to-pass rate @5. We refer to the five tests from Otter++ masks as planner, full, testLoc, patchLoc, and none. Planner represents the default Otter-generated tests where self-reflective action planning was applied. The other masks partially or completely expose localization information to the model to generate diverse solutions. For example, in ‘testLoc’, the test generation prompt contains only test localization outputs, but no patch localization or planner outputs.
Going beyond masks, this paper introduces another approach for heterogeneous prompting, adding five issue description morphs. Each morph rewrites the issue with an LLM, then runs Otter on the rewritten issues to generate diverse solutions. The goal is to increase the fail-to-pass rate @10. The five morphs are:
i) standard: This morph asks the LLM to standardize the issue description. A well-formatted issue description has several components, including a title, description of the problem, steps to reproduce, expected behavior, and actual behavior. Having these components may help the model to reason better or perform better in producing the bug reproduction test.
ii) simple: Some issue descriptions come with code artifacts and jargon that are very project-specific, making them hard to understand for a developer who does not work on the project regularly. We ask the LLM to simplify the issue so that it becomes easier to understand for developers.
iii) dropCode: We have seen some issue descriptions that come with misleading code snippets. Developers sometimes casually write snippets that may be useful for them, but are confusing to the model. For example, in Figure 3(a), the issue description uses Numpy (np) to manipulate a matrix. However, the Sympy project does not use or even install Numpy as a dependency for that particular issue. Seeing the original unmorphed issue, the model generates a test using np for matrix manipulation, which is not applicable in this setup. Dropping this code snippet helps the model avoid this mistake. The developer-written test in Figure 3(b) does not use Numpy.
iv) initTest: This morph asks the LLM to produce an initial fail-to-pass test and incorporate it into the issue description without giving any additional context. The rest of the pipeline may get some initial idea from the test and modify it so that it goes from fail to pass. Note that in this augmentation process, the model may contaminate the issue from a memorized test. We manually validate the possibility of contamination by the LLM and found that in 2% of cases, we observed some possible contamination where the model might have used memorized data to resolve the problem (see details in Section 5.1). Also, we are not performing anything better than zero-shot prompting here, and we observed that zero-shot prompting is not as effective as other approaches.
v) initPatch: Like initTest, this morph asks the model to propose an initial solution to resolve the issue, which may guide the rest of the pipeline to perform well. We found that in practice, for initTest and initPatch, the models sometimes do not write an initial test or patch due to a lack of confidence. The model may simply rewrite the issue with minor modifications.
We rerun the Otter pipeline using these new morphed issue descriptions and generate five new sets of tests. Note that issue rewriting may hurt performance since it may remove necessary information and add irrelevant context. However, our goal is to increase the $F\!\!\to\!\!P$ rate @ N, which could be achieved by such rewriting. Prompts and examples for execution-augmented test repair and issue description morphs are in the supplementary material.
3.2. Code Patch Generator
Our approach uses generated code patches to help with test selection. Prior work does something similar but in the other direction, using generated tests to help with code selection. CodeT (Chen et al., 2023) generates both code samples and corresponding test cases, then executes the generated code samples using the generated test cases and performs a dual execution agreement to choose the best code sample. Agentless (Xia et al., 2025) and later CodeMonkeys (Ehrlich et al., 2025) also use reproduction tests to select patches from multiple candidates but do not evaluate the quality of the reproduction test. Our e-Otter++ uses code patches to evaluate the quality of generated reproduction test candidates and attempt to choose the best one. We use Agentless patches, as opposed to patches from other systems, because their artifacts are open-source and their generated patches are available. Agentless generates a total of 40 code patches (1 greedy at temperature 0 and 39 non-greedy at temperature 0.8) prior to patch selection. We take the normalized version of these 40 patches and select the 5 most frequent ones. We chose 5 considering the number of existing normalized unique patches and our resource constraints.
3.3. Test Selection
The e-Otter++ test selector starts from 10 candidate refined tests and 5 candidate code patches, and its job is to pick 1 single final output test. The test selector is entirely rule-based and does not call any LLM. It starts by running each of the candidate tests both on $c_{\textrm{old}}$ and on each of the code patches $c_{\textrm{new}}$. If none of the tests are fail-to-pass on any code patch, then the test selector keeps all 10 candidate tests until the tie-breaking step, otherwise it:
  • Keeps only tests that are fail-to-pass.
  • Classifies those fail-to-pass tests into three groups based on how they failed on $c_{\textrm{old}}$:
    • assertion failure (these are most likely to have failed for the right reason),
    • other failure (the test ran but produced the wrong output), or
    • error (the test did not run properly, e.g., because of wrong syntax or an exception in a fixture; least likely to have failed for the right reason).
  • Keeps only the tests in the first non-empty of those three groups.
At this point, the test selector uses coverage as a tie-breaker for the remaining tests, by picking one test with the highest coverage. The coverage of a test $y$ is the average of its coverage on the 5 code patches. The coverage of a test $y$ on a code patch $c_{\textrm{new}}$ is the sum of the number of covered lines deleted from $c_{\textrm{old}}$ plus the number of covered lines added by $c_{\textrm{new}}$, divided by the total added+deleted lines. We refer to the test selected at this phase as e-Otter++ test.
4. Evaluation
This section describes experiments to answer the following research questions:
  • RQ1: How does our approach perform at generating $F\!\!\to\!\!P$ tests?
  • RQ2: How effective is execution-augmented test repair?
  • RQ3: How do issue description morphs and test-generation context masks increase the $F\!\!\to\!\!P$ @ N?
  • RQ4: How do generated code patches affect test selection?
  • RQ5: How does our approach help in validating code patches?
We evaluate our approach on two benchmarks: TDD-Bench Verified [^2] https://github.com/IBM/TDD-Bench-Verified and SWT-bench Lite [^3] https://github.com/logic-star-ai/swt-bench. Both benchmarks are derived from the popular issue resolution benchmark, SWE-bench (Jimenez et al., 2024). We report the fail-to-pass ($F\!\!\to\!\!P$) rate to show the effectiveness of the approach following prior works (Mündler et al., 2024; Nashid et al., 2025; Ahmed et al., 2025). We have primarily used two models, GPT-4o and Claude-3.7-sonnet. To have a fair comparison with Otter and Otter++, we used the same version of the GPT-4o model (gpt-4o-2024-08-06) as in their paper (Ahmed et al., 2025).
4.1. Performance of e-Otter and e-Otter++ (RQ1)
Table 1. Performance of e-Otter, e-Otter++, and baselines
Benchmark Approach $F\!\!\to\!\!P$ In (%)
TDD-Bench
Verified (449
samples)
Zero-shot (GPT-4o) 84 18.7
Otter++ (GPT 4o) 166 37.0
Otter++ (Claude-3.7-Sonnet) 174 38.8
e-Otter (GPT-4o) 167 37.2
e-Otter (Claude-3.7-Sonnet) 179 39.9
e-Otter++ (GPT-4o) 231 51.4
e-Otter++ (Claude-3.7-Sonnet) 283 63.0
SWT-bench
Lite (276
samples)
ZeroShot (GPT-4) 16 5.8
LIBRO (GPT-4) 42 15.2
AutoCodeRover (GPT-4) 25 9.1
SWE-Agent+ (GPT-4) 53 19.2
Otter++ (GPT 4o) 80 29.0
Issue2Test (GPT-4o-mini) 84 30.4
Otter++ (Claude-3.7-Sonnet) 84 30.4
Amazon Q Developer Agent* (Unknown) 104 37.7
e-Otter (GPT-4o) 80 29.0
e-Otter (Claude-3.7-Sonnet) 99 36.0
e-Otter++ (GPT-4o) 111 40.2
e-Otter++ (Claude-3.7-Sonnet) 145 52.5
*Number is taken from SWT-bench leaderboard.
Table 1 compares e-Otter and e-Otter++ with prior approaches on both TDD-Bench Verified and SWT-bench Lite. For TDD-Bench Verified and SWT-bench Lite, the best baselines report 38.8% and 37.7%, respectively. Our Claude-3.7-Sonnet based e-Otter++ achieves state-of-the-art performance on both benchmarks, achieving 63.0% and 52.5% on TDD-Bench Verified and SWT-bench Lite, respectively. We perform McNemar’s test (McNemar, 1947) between our best test generator and all prior approaches and it significantly outperforms all the baselines on both TDD-Bench Verified and SWT-bench Lite with a 99% confidence interval ($p<0.01$). McNemar’s test compares the performance of two systems on the same set of binary-labeled instances, highlighting statistically significant differences in their error patterns. We also report our performance with GPT-4 and GPT-4o to demonstrate model agnosticity. We find that the GPT-4o-based e-Otter++ significantly outperforms all prior GPT-4 and GPT-4o-based solutions. Table 1 omits some relatively weaker approaches from both benchmarks to reduce clutter. Apart from the works listed in Table 1, the SWT-bench Leaderboard has another entry, AEGIS, but hides it by default unless a “script mode” toggle is selected. AEGIS tests are not integrated with the project’s testing framework, unlike other leaderboard entries and e-Otter++. Under that setting, AEGIS reports a 47.8% $F\!\!\to\!\!P$ rate with an undisclosed model. Integrating tests with a testing framework enables future regression testing and coverage measurements. Although not directly comparable to AEGIS, e-Otter++ with Claude-3.7-Sonnet outperforms AEGIS on SWT-bench Lite with a 52.5% $F\!\!\to\!\!P$ rate. Note that, as mentioned in Section 2, the two benchmarks have different implementations. We primarily used the TDD-Bench harness because of its speed. We also ran the SWT-Bench harness and observed a slight drop in performance (50.7%) due to dependency issues and regression tests.
While the SWT-bench paper (Mündler et al., 2024) only describes the Lite subset of their benchmark, the leaderboard recently also added a Verified subset. SWT-bench Verified is similar to TDD-Bench Verified, as both are derived from SWE-Bench-Verified (Chowdhury et al., 2024). Although both benchmarks have some non-overlapping samples, systems usually perform similarly on them. For example, Otter++ (GPT-4o) reports 37.0% on both TDD-Bench Verified and SWT-bench Verified. Given this similarity between the two “verified” benchmarks, we opted to use only one of them to discuss our results in this paper.
Finding 1. e-Otter++ (Claude-3.7-Sonnet) significantly outperforms all prior approaches on TDD-Bench Verified and SWT-bench Lite, achieving 63% and 52.5% $F\!\!\to\!\!P$ rates.
4.2. Effectiveness of Test Repair (RQ2)
Table 2. Effectiveness of execution-augmented test repair
Benchmark Prompt Claude-3.7-Sonnet GPT-4o
Otter $F\!\!\to\!\!P$ e-Otter $F\!\!\to\!\!P$ Change Otter $F\!\!\to\!\!P$ e-Otter $F\!\!\to\!\!P$ Change
TDD-Bench
Verified (449
samples)
Planner 145 (32.3%) 179 (39.9%) 23.4% 141 (31.4%) 167 (37.2%) 18.4%
Context
Masking
full 125 (27.8%) 147 (32.7%) 17.6% 110 (24.5%) 135 (30.1%) 22.7%
testLoc 117 (26.1%) 130 (29.0%) 11.1% 115 (25.6%) 135 (30.1%) 17.4%
patchLoc 105 (23.4%) 134 (29.8%) 27.6% 107 (23.8%) 126 (28.1%) 17.8%
none 100 (22.3%) 127 (28.3%) 27.0% 110 (24.5%) 137 (30.5%) 24.5%
Issue
Morphing
standard 137 (30.5%) 183 (40.8%) 33.6% 104 (23.2%) 135 (30.1%) 29.8%
simple 137 (30.5%) 182 (40.5%) 32.8% 104 (23.2%) 128 (28.5%) 23.1%
dropcode 121 (26.9%) 168 (37.4%) 38.8% 77 (17.1%) 113 (25.2%) 46.8%
initTests 137 (30.5%) 166 (37.0%) 21.2% 101 (22.5%) 139 (31.0%) 37.6%
initPatch 151 (33.6%) 191 (42.5%) 26.5% 119 (26.5%) 149 (33.2%) 25.2%
SWT-bench
Lite (276
samples)
Planner 83 (30.1%) 99 (35.9%) 19.3% 70 (25.4%) 80 (29.0%) 14.3%
Context
Masking
full 67 (24.3%) 89 (32.2%) 32.8% 58 (21.0%) 71 (25.7%) 22.4%
testLoc 67 (24.3%) 82 (29.7%) 22.4% 55 (19.9%) 76 (27.5%) 38.2%
patchLoc 59 (21.4%) 72 (26.1%) 22.0% 53 (19.2%) 62 (22.5%) 17.0%
none 57 (20.7%) 68 (24.6%) 19.3% 58 (21.0%) 75 (27.2%) 29.3%
Issue
Morphing
standard 80 (29.0%) 95 (34.4%) 18.8% 62 (22.5%) 76 (27.5%) 22.6%
simple 68 (24.6%) 88 (31.9%) 29.4% 57 (20.7%) 69 (25.0%) 21.1%
dropcode 72 (26.1%) 94 (34.1%) 30.6% 35 (12.7%) 61 (22.1%) 74.3%
initTests 77 (27.9%) 99 (35.9%) 28.6% 59 (21.4%) 68 (24.6%) 15.3%
initPatch 73 (26.4%) 92 (33.3%) 26.0% 68 (24.6%) 76 (27.5%) 11.8%
Table 2 shows the $F\!\!\to\!\!P$ rate for each of the 10 tests before (Otter) and after (e-Otter) execution-augmented test repair. For example, on TDD-Bench Verified, starting from the default Planner test, using Claude-3.7-Sonnet to repair tests improves performance from 32.3% to 39.9% $F\!\!\to\!\!P$ rate, a relative change of 23.4%. While the improvements vary by test, model, and benchmark, they are strong across the board, with relative changes ranging from a minimum of 11.1% to a maximum of 74.3%.
Contributions of Different Components of Execution-augmented Test Repair.
Table 3 shows the contribution of each component of execution-augmented test repair. Due to cost and resource constraints, we performed this experiment with the GPT-4o model on TDD-Bench Verified only. The row labeled e-Otter shows the performance starting with the Planner test and all repair components intact. The remaining rows show the performance when individual components are selectively disabled. The results show that each component contributes 7.8% to 21.0% to e-Otter. To see the effectiveness of the critic, we remove the critic and fix the problem only once instead of making several attempts because the subsequent attempts depend on the critic. For temperature evaluation, we simply replace 0.8 with 0, i.e., greedy. Note that the critic plays the most important role in test repair.
Table 3. Contributions of different components of execution-augmented test repair (ablation)
Approach $F\!\!\to\!\!P$ Change
e-Otter 167 (37.2%) -
without Buggy Line Selection 152 (33.9%) -9.0%
without Relevant Line Extraction 154 (34.3%) -7.8%
without Function Lookup 153 (34.1%) -8.4%
without Critic 132 (29.4%) -21.0%
without High Temperature (0.8) 155 (34.5%) -7.2%
Table 4. LLMs as critic in test repair
Approach TP FP FN Precision Recall F-Score
Golden Test 321 0 128 1.00 0.71 0.83
Otter 127 49 14 0.72 0.90 0.80
Otter++ 151 68 15 0.69 0.91 0.78
How good are LLMs as critics?
We just saw that that critics play an important role in test repair. But how good are LLMs as critics? We use a GPT-4o based critic on different tests and see how it performs. We take GPT-4o based Otter, Otter++, and golden (i.e., developer-written) tests from TDD-Bench Verified. We feed the test and test log and ask the model to comment on whether the test is failing for the right reason mentioned in the issue description. Table 4 shows that GPT-4o achieves a 0.78 to 0.83 F-score as a critic, which indicates the benefit of using LLMs as critics.
Figure 4. Number of attempts taken in test repair
How many attempts are required to repair the test?
The execution-augmented test repair loop iterates for at most 10 attempts. We measured the actual number of attempts on the 449 instances of TDD-Bench Verified with all 5 morphs + 5 masks, i.e., 4,490 tests. Figure 4 shows the highest sample distributions at 0 or 10 attempts. There are many tests which are already $F\!\!\to\!\!P$ and do not need any repair. If the critic identifies this situation correctly, there are zero repair attempts. For many tests, even after 10 attempts, the critic is not satisfied or is only satisfied at the $10^{\textrm{th}}$ attempt, and those are presented by “10” in the plot. The $F\!\!\to\!\!P$ rate is at least 10% if the critic stops at 8 attempts. We can certainly go beyond 10 attempts, but after 8 attempts, the $F\!\!\to\!\!P$ rate drops significantly and is less than 2% at 10 attempts.
Finding 2. e-Otter significantly improves the $F\!\!\to\!\!P$ rate. The improvement ranges from 11.1% to 74.3% based on the model and benchmark under consideration.
4.3. Impact of Heterogeneous Prompts (RQ3)
All $F\!\!\to\!\!P$ numbers we reported so far were “@ 1”, measuring the rate at which a single test is fail-to-pass. Here, we instead explore $F\!\!\to\!\!P$ @ N, i.e., the rate at which at least one among N generated tests is fail-to-pass. That is an upper bound for how well a system can do by picking a single test to submit from multiple generated candidate tests. Morphs and masks yield heterogeneous prompts, which in turn increase the diversity of candidate tests, thus hopefully increasing $F\!\!\to\!\!P$ @ N. We reuse 5 heterogeneous prompts including masks from Otter++ (Ahmed et al., 2025) and add 5 more prompts using morphs for a total of 10 heterogeneous prompts. Table 5 shows that morphs increase $F\!\!\to\!\!P$ @ N, even without masks, both before (Otter) and after (e-Otter) test repair. On e-Otter, the Claude-3.7-Sonnet model with “plan + mask + morph” increases $F\!\!\to\!\!P$ @ N by 17.1% and 23.8% compared to “plan + mask” on SWT-bench Lite and TDD-Bench Verified, respectively. We observe similar improvements with the GPT-4o model on both benchmarks.
Table 5. Impact of heterogeneous prompts (morphs & masks)
Benchmark Model Category N Otter $F\!\!\to\!\!P$ @ N Change e-Otter $F\!\!\to\!\!P$ @ N Change
TDD-Bench
Verified
Claude-3.7-
Sonnet
plan + mask 5 210 (46.8%) - 260 (57.9%) -
plan + morph 6 255 (56.8%) 21.4% 305 (67.9%) 17.3%
plan + mask + morph 10 270 (60.1%) 28.6% 322 (71.7%) 23.8%
GPT-4o plan + mask 5 197 (43.9%) - 219 (48.8%) -
plan + morph 6 211 (47.0%) 7.1% 240 (53.5%) 9.6%
plan + mask + morph 10 237 (52.8%) 20.3% 260 (57.9%) 18.7%
SWT-bench
Lite
Claude-3.7-
Sonnet
plan + mask 5 113 (40.9%) - 146 (52.9%) -
plan + morph 6 141 (51.1%) 24.8% 162 (58.7%) 11.0%
plan + mask + morph 10 147 (53.3%) 30.1% 171 (62.0%) 17.1%
GPT-4o plan + mask 5 105 (38.0%) - 119 (43.1%) -
plan + morph 6 114 (41.3%) 8.6% 126 (45.7%) 5.9%
plan + mask + morph 10 124 (44.9%) 18.1% 135 (48.9%) 13.4%
Figure 5. $F\!\!\to\!\!P$ @ N vs. N: a comparison of tests generated with higher temperature and heterogeneous prompting.
Why not use high temperature instead of heterogeneous prompting?
Prior works (Xia et al., 2025; Ehrlich et al., 2025; Chen et al., 2021) mostly depend on high temperature for multi-sampling. Figure 5 compares the results of heterogeneous prompting with the samples generated with high temperature. We tried three distinct temperatures (0.5, 0.8, and 1.0) and find that the best results are obtained at a temperature of 0.8, which is consistent with prior research (Chen et al., 2021). However, our approach, “plan + mask + morph”, significantly outperforms samples generated with high temperature. In Figure 5, for each $N$ (1-10), we take all possible combinations from 10 candidates and show the confidence interval in the plot.
Table 6. Contribution of each prompt to $F\!\!\to\!\!P$ @ N
Approach Claude-3.7-Sonnet GPT-4o
$F\!\!\to\!\!P$ Change $F\!\!\to\!\!P$ Change
All (plan + mask + morph) 322 (71.7%) - 260 (57.9%) -
without planner 316 (70.4%) -1.9% 248 (55.2%) -4.6%
without full 317 (70.6%) -1.6% 259 (57.7%) -0.4%
without testLoc 318 (70.8%) -1.2% 258 (57.5%) -0.8%
without patchLoc 321 (71.5%) -0.3% 259 (57.7%) -0.4%
without none 319 (71.0%) -0.9% 251 (55.9%) -3.5%
without standard 318 (70.8%) -1.2% 256 (57.0%) -1.5%
without simple 315 (70.2%) -2.2% 259 (57.7%) -0.4%
without dropCode 318 (70.8%) -1.2% 253 (56.3%) -2.7%
without initTest 317 (70.6%) -1.6% 254 (56.6%) -2.3%
without initPatch 312 (69.5%) -3.1% 256 (57.0%) -1.5%
Impact of each prompt in $F\!\!\to\!\!P$ @ N (ablation)
Table 6 shows the impact of each prompt for the Claude-3.7-Sonnet and GPT-4o models on TDD-Bench Verified. Each prompt contributes between 0.3% and 4.6% to $F\!\!\to\!\!P$ @ N. Even though some prompts do not add many new instances, increasing $F\!\!\to\!\!P$ @ N is still beneficial, as if gives the test selector a larger pool of valid candidates to choose from, increasing the chances of selecting the most suitable ones.
Although we report most results using two closed-source models, Claude-3.7-Sonnet and GPT-4o, we also investigate the effectiveness of execution-augmented test repair and heterogeneous prompting on a relatively weaker open-source model, Mistral-large (123 billion parameters). The $F\!\!\to\!\!P$ rate increases by 8.3% to 37% by execution-augmented test repair, and $F\!\!\to\!\!P$ @ N increases by 11% with heterogeneous prompting (plan + mask + morph). This indicates that our approaches are also applicable to open-source models.
Finding 3. The samples with “plan + mask + morph” increase $F\!\!\to\!\!P$ @ N by more than 13% compared to “plan + mask”, regardless of the model and benchmark.
4.4. Effectiveness of Code Patches (RQ4)
Table 7. Test selection using code patches
Benchmark Model Pre-filter Selection $F\!\!\to\!\!P$
TDD-Bench
Verified
Clade-3.7-Sonnet None Random 200.5 (44.7%)
None Coverage 213.0 (47.4%)
CodePatch Random 277.6 (61.8%)
CodePatch Coverage 283.0 (63.0%)
GPT-4o None Random 161.6 (36.0%)
None Coverage 182.0 (40.5%)
CodePatch Random 227.0 (50.6%)
CodePatch Coverage 231.0 (51.4%)
SWT-bench
Verified
Clade-3.7-Sonnet None Random 100.9 (36.5%)
None Coverage 111.0 (40.2%)
CodePatch Random 140.6 (50.9%)
CodePatch Coverage 145.0 (52.5%)
GPT-4o None Random 83.4 (30.2%)
None Coverage 88.0 (31.9%)
CodePatch Random 108.4 (39.3%)
CodePatch Coverage 111.0 (40.2%)
This section examines the effectiveness of test selection using execution feedback from generated code patches to pre-filter a test patch, as well as using coverage as a final selection tie-breaker. The baseline for the tie-breaker is random selection, where we repeat the experiment 10 times and report the average $F\!\!\to\!\!P$ rate. Table 7 shows the results for both models on both benchmarks. Without code patches as a pre-filter, coverage alone improves performance by 6.2% (from 200.5 to 213) for the Claude model on TDD-Bench Verified. With generated code patches as a pre-filter, coverage still provides some benefit, but the improvement is relatively low, at 1.9% (from 277.6 to 283). In contrast, code patches have a significant impact on performance. With coverage as a tie-breaker, the $F\!\!\to\!\!P$ rate increases by 32.9% (from 213 to 283) because of pre-filtering. Other model-benchmark combinations show similar results.
Finding 4. Test selection benefits from both the use of generated code patches as a pre-filter and coverage as a tie-breaker among multiple tests within a group. Generated code patches play the most important role in test selection.
4.5. Validating SWE-patches (RQ5)
Table 8. Validating SWE-patches with our generated tests
Rank System Old Precision Claude-3.7-Sonnet GPT-4o
Recall New Precision Recall New Precision
1 tools_claude-4-opus 74.2% 72.6% 78.8% 57.7% 81.5%
2 tools_claude-4-sonnet 73.5% 73.3% 78.1% 59.0% 79.6%
9 sweagent_claude-4-sonnet 68.4% 69.9% 74.5% 56.6% 74.8%
10 aime_coder 67.9% 58.8% 78.4% 47.9% 82.3%
11 openhands 66.1% 72.8% 73.4% 58.8% 75.4%
The tests generated by our approach can be used for discarding bad code patches and increasing the precision of solutions proposed by different systems from the SWE-bench Verified leaderboard. Ahmed et al. (Ahmed et al., 2025) took the top 22 systems from the leaderboard and ran the five tests generated by Otter++. They filtered out a code patch if all the tests failed on it. This achieved a precision of 65% to 92% while maintaining a decent recall of 30%-41%. Mündler et al. (Mündler et al., 2024) performed a similar experiment with one system. Note that neither of these two prior approaches could achieve a $F\!\!\to\!\!P$ rate beyond 50% even with 5 samples. Therefore, the recall suffers significantly. In contrast, e-Otter++ achieves a 71.7% $F\!\!\to\!\!P$ @ 10 rate, which could significantly benefit the recall. To showcase the efficacy of our tests in conjunction with top SWE agents, we pick the top SWE agents from the leaderboard where the generated code patches were available. Unfortunately, due to a recent change in the SWE-Bench submission process, we could not retrieve them for some systems. Table 8 shows the old precision, recall, and new precision on SWE-Bench-Verified using both Claude-3.7-Sonnet and GPT-4o models. We achieve 47.9% to 72.8% recall, regardless of the model, while substantially improving precision.
Finding 5. We achieve 47.9% to 72.8% recall, regardless of the model, while increasing the precision of the SWE-patches.
5. Discussion
This section briefly discusses model contamination, test coverage, and a detailed analysis of the tests.
5.1. Contamination through Issue Morphs
Test generation from issues could potentially suffer from model memorization and contamination since the model might have seen the data during the training phase. Both prior works by Ahmed et al. (Ahmed et al., 2025) and Mündler et al. (Mündler et al., 2024) addressed this issue and show that the model-generated tests are very different from developer-written tests and the impact of contamination is minimal. Since we build upon Ahmed et al.’s Otter++, we refer to their contamination analysis except for places where e-Otter++ might potentially introduce new contamination. Specifically, issue morphs might introduce contamination while rewriting the issue. To investigate this, we manually validated the samples before execution-augmented test repair. We took 50 samples (25 from each model) where the original planner prompt was not producing $F\!\!\to\!\!P$ tests but after issue morph the tests became $F\!\!\to\!\!P$. We had 4 raters with programming experience ranging from 15-30 years. We presented the raters with four pieces of information: (i) original issue description, (ii) modified issue, (iii) non $F\!\!\to\!\!P$ version of the test, (iv) $F\!\!\to\!\!P$ version of the test. We asked the raters the following three questions.
  • Is the modified issue adding new information that might influence the model’s decision?
  • Is the modified issue removing any information that might influence the model’s decision?
  • In cases of addition, does this rewrite cause contamination that might help the model generate fail-to-pass tests? The rater can decide on this by comparing the two versions of the tests.
The raters initially rated 5 samples and had a discussion on how to rate the remaining samples. After that, each rater individually rated the samples. Since we have four raters, it is infeasible to do majority voting on the rated instances. Instead, we adapted a stricter metric to assign the final label. If at least two raters think there is some addition, removal, or contamination, we counted it as “1”, otherwise “0”.
Table 9 shows the inter-rater agreement and outcome of the study. We use Fleiss’ Kappa (Fleiss, 1971) to compute inter-rater agreement, and the raters show fair to substantial agreement [^4] Kappa ¡ 0: Poor, 0.01–0.20: Slight, 0.21–0.40: Fair, 0.41–0.60: Moderate, 0.61–0.80: Substantial, 0.81–1.00: Almost perfect agreement.. Note that for contamination, we observe a very low kappa value because all raters mostly choose the same category. This phenomenon is known as the kappa paradox (Feinstein and Cicchetti, 1990). Overall, 40% of modified issues have some information addition and 68% of issues have some information removal. Claude-3.7-Sonnet makes more additions and removals compared to GPT-4o. However, none of the raters think there is substantial contamination due to issue morphs. Only for 1 (2%) sample out of 50 samples, more than one rater think the rewriting might have caused some contamination. Therefore, we can infer that issue morph does not contaminate the issue. We share the ratings from each individual rater in the supplementary material.
Table 9. Manual validation to investigate model contamination or memorization
Category Model N Fleiss’ Kappa Count
Value Interpretation
Addition Claude-3.7-Sonnet 25 0.33 Fair 16 (64%)
GPT-4o 25 0.37 Fair 4 (16%)
Both 50 0.44 Moderate 20 (40%)
Removal Claude-3.7-Sonnet 25 0.57 Moderate 19 (76%)
GPT-4o 25 0.63 Substantial 15 (30%)
Both 50 0.60 Moderate 34 (68%)
Contamination Claude-3.7-Sonnet 25
Kappa value is not meaningful
because all raters mostly choose
the same category.
0 (0%)
GPT-4o 25 1 (4%)
Both 50 1 (2%)
Figure 6. Example initTest (sympy__sympy-14711).
Figure 6 shows the example that was marked by three out of four raters as possible contamination. When we took a closer look at it, we became uncertain about it. In this particular case, the LLM was asked to generate an initial test with the issue description. The fixed line is “result = sum([N.x, (0 * N.x)])”. This line is present in the modified issue description and that might have inspired the raters to mark it as a suspicious test. However, that same line was also present in the original issue description. The original test considered that line buggy and opted for a different version of this line. Seeing the fixed line in the proposed test definitely helps the model to use the right version. But whether this case constitutes actual contamination is unclear since that line is present in the original issue description.
5.2. Coverage of Generated Tests
Coverage is an important metric to measure the adequacy of tests. We measure how well a generated test covers the developer-written golden code patch (note that this is not available during test generation). To do that, we find the deleted lines in $c_{\textrm{old}}$ and added lines in $c_{\textrm{new}}$ and see whether the test covers those lines. We take the total number of covered lines and divide them by the total number of deleted and added lines to compute the coverage. Figure 7 shows that $F\!\!\to\!\!P$ tests have higher coverage compared to other tests. These findings are consistent with prior works (Ahmed et al., 2025; Mündler et al., 2024) and motivate using coverage as a tie-breaker for test selection (of course, test selection uses coverage on generated code patches, not golden ones).
Figure 7. Coverage of the generated tests
5.3. Detailed Analysis of the Tests
Table 10 presents an analysis of the e-Otter-generated tests using GPT-4o from different perspectives. We did not see much impact of focal localization on the performance, with 37.8% and 34.2% for correct and incorrect localization, respectively. Test localization has a significant impact on the performance (42.6% vs. 24.2% for correct and incorrect localization). We see that the fail-to-pass rate is higher when modify an existing test (40.2%) than when writing a new test (36.1%), which is not surprising because the model can have better context from the existing test. It is expected that the test will fail on the old codebase. However, we found 43 samples in e-Otter where the test passed on $c_{\textrm{old}}$. Also, our analysis shows that tests with assertion failures have a higher success rate (50.4% fail-to-pass rate) compared to other groups.
Table 10. Analysis of the e-Otter-generated tests using GPT-4o model from different perspectives.
Perspective Category #of Sample $F\!\!\to\!\!P$ In (%)
Focal Localization Correct 370 140 37.8
Wrong 79 27 34.2
Test Localization Correct 317 135 42.6
Wrong 132 32 24.2
Type of Test Modified 122 49 40.2
New 327 118 36.1
Test on old Codebase Pass 43 0 0
AssertionFail 224 113 50.4
Fail 94 46 48.9
Error 88 8 9.1
6. Threats to Validity
One limitation of our work is that it only experiments with Python. As a result, the findings may not generalize to other programming languages or repositories outside the benchmarks. However, it is worth noting that prior works (Jimenez et al., 2024; Xia et al., 2025; Zhang et al., 2024), which share this same limitation, have still been impactful and contributed significantly to the field.
One of the key concerns with SWE-bench related work is the contamination or memorization problem of the model. The model may have seen the data during the training process, and that might positively impact the results. Prior works (Jimenez et al., 2024; Ahmed et al., 2025; Mündler et al., 2024) have attempted to address this problem. The SWE-bench paper claims that “difficulty does not correlate with issue resolution date”, indicating minimal contamination. Ahmed et al. show that Otter-generated tests differ substantially from developer-generated tests. In this work, we reuse Otter-generated tests. The only way we can introduce new contamination is while rewriting the issue description for morphs. However, Section 5.1 demonstrates through manual validation that issue morphs do not introduce contamination. We believe our results are not influenced by major model contamination.
Our system, e-Otter++, uses both sequential inference scaling (via test repair) and parallel inference scaling (via heterogeneous prompting). This increases the number of LLM calls. Furthermore, we also use Agentless code patches for test selection, which requires several LLM calls. We did not generate the code patches for Agentless ourselves; instead, we relied on code patches shared by the authors for both the Claude and GPT-4o models. However, based on their reported numbers and our experiments, we can estimate the cost for each sample. For the Claude and GPT-4o models, our estimated cost is \$2.75 and \$1.80 per instance, respectively. Note that prior test generation approaches do not use code patches for test selection, which adds more cost. However, it is likely that a reproduction test generation approach will be paired with a SWE-patch generation approach. In that case, the additional cost of generating code patches will be amortized.
Another limitation of e-Otter++ is that it considers only one test file and generates only one block of code. In real-world projects, test code can be spread across multiple files or blocks of code. Despite that, e-Otter++ exceeds the state-of-the-art performance, so we leave further improvements to future work.
7. Related Work
There has been a flurry of recent papers about generating reproduction tests from issues, which is the same problem we tackle in our paper. Libro generates tests for Java bug reports (Kang et al., 2023). It repairs tests based on compiler feedback, not execution feedback like we do, and generates several tests using high LLM decoding temperature, not morphs and masks like we do. Plein et al. also explore the feasibility of generating tests for Java bug reports, albeit without test repair (Plein et al., 2024). Mündler et al. (Mündler et al., 2024) introduce SWE-Agent+, a Python reproduction test generation agent derived from the SWE-Agent (Yang et al., 2024) code patch generator. While SWE-Agent+ has access to tools that can give it execution feedback, it leverages such feedback less effectively than e-Otter++. The same holds for EvoCoder (Lin et al., 2024), also derived from SWE-Agent, and for USEAgent (Applis et al., 2025), derived from AutoCodeRover (Zhang et al., 2024). AEGIS is a novel multi-agent system for reproduction test generation (Wang et al., 2024). It repairs tests based on execution feedback from $c_{\textrm{old}}$, but unlike our work, does not augment the test repair prompt. Also, it standardizes issue descriptions, but unlike our work, does not use morphs or masks to increase test candidate pool diversity. Otter++ (Ahmed et al., 2025) is an LLM-based workflow for reproduction test generation that we use as a component in e-Otter++, which extends it in three important ways: (i) augmenting the test repair prompt; (ii) using morphs (Otter++ only uses masks); and (iii) leveraging execution feedback from generated candidate code patches. Our experiments show that these novel contributions greatly improve the $F\!\!\to\!\!P$ rate. BRT Agent is a new agent for generating reproduction tests (Cheng et al., 2025), and Issue2Test is a new non-agentic workflow for the same task (Nashid et al., 2025); neither augments test repair prompts, uses masks or morphs, or leverages candidate code patches.
Unlike e-Otter++ and the above papers on using generated code patches to help pick or improve reproduction tests, there is also work on the opposite: using reproduction tests to help pick or improve code patches. Papers on using tests to help pick (but not improve) code patches include Smith et al.’s 2015 paper (Smith et al., 2015), CodeT (Chen et al., 2023), Agentless (Xia et al., 2025), and R2E-Gym (Jain et al., 2025). Similarly, several papers use tests to help improve (but not pick) code patches (Zhang et al., 2024; Yang et al., 2024; Chen et al., 2024; Arora et al., 2024; Ruan et al., 2024; Mathews and Nagappan, 2024). Some papers even do both, using tests to help pick and improve code patches: CodeMonkeys (Ehrlich et al., 2025) runs several parallel workflows, each of which iteratively refines a $\langle c_{\mathrm{new}},y\rangle$ pair; and PatchPilot (Li et al., 2025) uses reproduction tests for bug localization before code patch generation, and later uses them again to pick among multiple candidate code patches. However, while most of the work described in this paragraph uses generated tests, test generation is not their main focus: their reproduction test generation components are more basic, and they lack empirical evaluation of the $F\!\!\to\!\!P$ rate of generated tests. That said, they demonstrate a widespread need for our work.
8. Conclusion
This paper introduces e-Otter++, a system for generating reproduction tests from SWE issues. It works by first generating a set of candidate tests, with novel techniques for increasing diversity by morphing issue descriptions and masking parts of the test generation context. Next, it iteratively repairs each test using feedback from executing it on the old code; what is novel in this step is that it augments the test repair prompt with new context based on the execution feedback. Finally, it selects a single final reproduction test; since the new code that fixes the issue does not yet exist, it uses execution feedback from multiple generated candidate code patches instead to inform test selection. Through extensive experiments, this paper demonstrates that e-Otter++ outperforms prior work on this problem, and that these improvements can be attributed to the novel features of e-Otter++. The experiments also demonstrate that the tests generated by e-Otter++ are effective at boosting precision of SWE agents that generate patches from issues. Overall, we hope that by automatically generating reproduction tests, we can take the tedium out of test-driven development (TDD), so more developers and software projects can reap the benefits of TDD. All our generated tests and logs are available at https://doi.org/10.5281/zenodo.15860535.

### v2

@@
-Execution-Feedback Driven Test Generation from SWE Issues
+Heterogeneous Prompting and Execution Feedback for SWE Issue Test Generation and Selection
 Abstract.
@@
 LLMs, SWE Patches, Reproduction Tests
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
+journalyear: 2026copyright: rightsretainedconference: 2026 IEEE/ACM 46th International Conference on Software Engineering; April 15–17, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06
 1. Introduction
@@
 But such tests, when they exist, are invaluable for software development.
-Before issue resolution, tests can reproduce the issue to confirm whether it is real, make it more concrete, and help stakeholders agree upon the desired behavior—following the highly regarded software development practice of test-driven development (TDD) (Beck, 2002).
+Before issue resolution, tests can reproduce the issue to confirm whether it is real, make it more concrete, and help stakeholders agree upon the desired behavior—following the highly regarded software development practice of test-driven development (TDD) (beck_2002).
 Once a code change (a patch) is proposed to resolve the issue, running the same reproduction tests helps decide whether to accept the patch.
@@
 Unfortunately, generating $F\!\!\to\!\!P$ tests is challenging, because issue descriptions are informal and because the new code patch on which the tests should pass does not yet exist.
-In addition to helping human software engineers, generating tests from SWE issues can also assist automated SWE agents, as several SWE agents also rely on reproduction tests (Arora et al., 2024; Chen et al., 2024; Ehrlich et al., 2025; Jain et al., 2025; Li et al., 2025; Ruan et al., 2024; Xia et al., 2025; Yang et al., 2024; Zhang et al., 2024).
+In addition to helping human software engineers, generating tests from SWE issues can also assist automated SWE agents, as several SWE agents also rely on reproduction tests (arora_et_al_2024; chen_et_al_2024; ehrlich_et_al_2025; jain_et_al_2025; li_et_al_2025; ruan_zhang_roychoudhury_2024; xia_et_al_2025; yang_et_al_2024; zhang_et_al_2024).
 There have been several recent attempts at using large language models (LLMs) to automatically generate $F\!\!\to\!\!P$ tests from SWE issues, without access to the resolving code patches.
-Libro uses an LLM to directly generate 50 tests from Java bug reports, then performs rule-based postprocessing and ranking (Kang et al., 2023).
-Other attempts focus on Python, modifying a SWE agent designed to automatically generate code patches to generate tests instead (Lin et al., 2024; Mündler et al., 2024).
-Recent papers proposed novel LLM-based flows or agents custom-tailored to the problem of generating tests from issues, including AEGIS (Wang et al., 2024), Otter++ (Ahmed et al., 2025), BRT agent (Cheng et al., 2025), and Issue2Test (Nashid et al., 2025).
-There are two benchmarks for generating tests from issues: SWT-bench (Mündler et al., 2024) (derived from SWE-bench (Jimenez et al., 2024)) and TDD-Bench Verified (Ahmed et al., 2024) (derived from SWE-bench Verified (Chowdhury et al., 2024)).
+Libro uses an LLM to directly generate 50 tests from Java bug reports, then performs rule-based postprocessing and ranking (kang_yoon_yoo_2023).
+Other attempts focus on Python, modifying a SWE agent designed to automatically generate code patches to generate tests instead (lin_et_al_2024; mundler_et_al_2024).
+Recent papers proposed novel LLM-based flows or agents custom-tailored to the problem of generating tests from issues, including AEGIS (wang_et_al_2024), Otter++ (ahmed_et_al_2025), BRT agent (cheng_et_al_2025), and Issue2Test (nashid_et_al_2025).
+Table 1 shows how e-Otter++ is procedurally distinct.
+Note that Otter, Issue2Test, and e-Otter do not use inference scaling or test selection.
+There are two benchmarks for generating tests from issues: SWT-bench (mundler_et_al_2024) (derived from SWE-bench (jimenez_et_al_2024)) and TDD-Bench Verified (ahmed_et_al_2024-tdd) (derived from SWE-bench Verified (chowdhury_et_al_2024)).
 SWT-bench Verified, also derived from SWE-bench Verified, is very similar to TDD-Bench Verified.
@@
 However, as there is no paper about Amazon Q Developer Agent, its inner workings are not publicly known.
-The top system on SWT-bench Verified with a published paper is Otter++ (Ahmed et al., 2025) with an $F\!\!\to\!\!P$ rate of 37.0%, and it also has the exact same $F\!\!\to\!\!P$ rate of 37.0% on TDD-Bench Verified.
+The top system on SWT-bench Verified with a published paper is Otter++ (ahmed_et_al_2025) with an $F\!\!\to\!\!P$ rate of 37.0%, and it also has the exact same $F\!\!\to\!\!P$ rate of 37.0% on TDD-Bench Verified.
 Given how similar the two verified benchmarks are, this paper evaluates on TDD-Bench Verified only.
@@
 The “e-” refers to execution feedback: e-Otter++ outperforms prior work by leveraging execution feedback more effectively.
+Additionally, we apply our approach to the contamination-free benchmark SWE-rebench (badertdinov2025swe) to show the effectiveness of our two primary contributions (execution feedback and heterogeneous prompting), as SWE-bench is suspected to be exposed to recent frontier models during training (liang2025swe).
 The first place where e-Otter++ uses execution feedback is after generating a candidate test.
@@
 Ultimately, we hope our work helps cement the practice of using tests both to drive issue resolution and for long-term continuous quality assurance after.
+Table 1. Component-wise comparison with prior works
+System Test Generation Inference Scaling Test Selection
+Execution
+Feedback
+Context
+Augment
+Critic
+Restart with
+Temp 0
+Mask Morph
+Execution
+on $c_{\textrm{old}}$
+Execution on
+Surrogate Patch
+Coverage
+Otter No No No Not Applicable Not Applicable
+Otter++ No No No No Yes No Yes No No
+Isue2Test Yes No Yes Not Applicable Not Applicable
+AEGIS Yes No Yes Yes No No Yes No No
+e-Otter Yes Yes Yes Not Applicable Not Applicable
+e-Otter++ Yes Yes Yes No Yes Yes Yes Yes Yes
 2. Problem Statement and Background
-This section discusses the problem statement and briefly introduces prior work we build upon: Otter (Ahmed et al., 2025), Agentless (Xia et al., 2025), SWT-bench (Mündler et al., 2024), and TDD-Bench Verified (Ahmed et al., 2024).
+This section discusses the problem statement and briefly introduces prior work we build upon: Otter (ahmed_et_al_2025), Agentless (xia_et_al_2025), SWT-bench (mundler_et_al_2024), and TDD-Bench Verified (ahmed_et_al_2024-tdd).
 Figure 1.
@@
 2.2. Reproduction Test Generation: Otter
-This paper introduces e-Otter++, which extends an earlier system called Otter++ (Ahmed et al., 2025).
+This paper introduces e-Otter++, which extends an earlier system called Otter++ (ahmed_et_al_2025).
 Both e-Otter++ and Otter++ generate tests directly from issue descriptions and $c_{\textrm{old}}$.
@@
 One of the new techniques in e-Otter++ is how it uses execution feedback from automatically generated candidate code patches.
-Rather than implementing a new code patch generator, it reuses an existing one, Agentless (Xia et al., 2025), as a component.
+Rather than implementing a new code patch generator, it reuses an existing one, Agentless (xia_et_al_2025), as a component.
 Agentless implements a hierarchical bug localization process.
@@
 Our new system e-Otter++ uses only the first part of Agentless, without the test generation or patch selection of Agentless, as a component (see Section 3.2).
-While we could have adopted any code patch generator, we chose Agentless because it is open-source and candidate patches are publicly available [^1] https://github.com/OpenAutoCoder/Agentless/tree/main.
-2.4. Test Evaluation Benchmarks:
-TDD-Bench Verified and SWT-bench Lite
+While we could have adopted any code patch generator, we chose Agentless because it is open-source and candidate patches are publicly available (agentless_artifact).
+2.4. Test Evaluation Benchmarks
 There are two benchmarks for evaluating bug reproduction tests.
-SWT-bench Lite (Mündler et al., 2024) was derived primarily from the SWE-bench Lite dataset (Jimenez et al., 2024) and contains 276 samples.
-TDD-Bench Verified (Ahmed et al., 2024) was derived from SWE-bench Verified (Chowdhury et al., 2024).
+SWT-bench Lite (mundler_et_al_2024) was derived primarily from the SWE-bench Lite dataset (jimenez_et_al_2024) and contains 276 samples.
+TDD-Bench Verified (ahmed_et_al_2024-tdd) was derived from SWE-bench Verified (chowdhury_et_al_2024).
 In TDD-Bench Verified, the authors ran 500 samples from SWE-bench Verified, filtered out samples with zero test coverage or those violating the $F\!\!\to\!\!P$ property of the tests, and ended up with 449 samples.
@@
 It also increases reproduction test candidate diversity via issue description morphs and test-generation context masks.
+Input : issue description $d_{\textrm{issue}}$; old code $c_{\textrm{old}}$; candidate test $y$
+Output : refined candidate test
+1 repeat
+2  $\textit{executionFeedback}\leftarrow\textit{executeTestOnCode}(y,c_{\textrm{old}})$;
+3  $\textit{critique}\leftarrow\textit{LLM}(\textit{criticPrompt}(d_{\textrm{issue}},\textit{executionFeedback}))$;
+4  if not $\textit{failsForRightReason}(\textit{critique})$ then
+5     $\textit{reads}\leftarrow\textit{getActions}(\textit{critique})$;
+6     $\textit{extraContext}\leftarrow[\textit{executeAction}(a,c_{\textrm{old}})\textbf{ for }a\in\textit{reads}]$;
+7     $y\leftarrow\textit{LLM}(\textit{repairPrompt}(d_{\textrm{issue}},y,\textit{extraContext})$
+8  end if
+9
+10until $\textit{failsForRightReason}(\textit{critique})$ or too many attempts;
+return $y$
+Algorithm 1 Execution-augmented test repair
 Execution-augmented Test Repair.
-The test repair component starts by running an Otter-generated candidate test $y$ on $c_{old}$.
+As shown in Algorithm 1, the test repair component starts by running an Otter-generated candidate test $y$ on $c_{old}$.
 It is expected that the test should fail in this phase.
@@
 Chen et al.
-(Chen et al., 2025) reported that Python execution errors during the issue resolution phase correlate with lower resolution rates.
+(chen_ma_jiang_2025) reported that Python execution errors during the issue resolution phase correlate with lower resolution rates.
 They also identified the most prevalent errors—such as ModuleNotFoundError and TypeError—in the generated code, which tend to differ greatly from the reason mentioned in the issue description.
@@
 Then, the repair loop starts over from this new test as $y$, until the critic is satisfied or for a maximum of 10 iterations.
-Following prior works (Xia et al., 2025; Ehrlich et al., 2025; Chen et al., 2021), the test repair LLM call uses a higher temperature (0.8) to encourage the model to try diverse solutions.
+Following prior works (xia_et_al_2025; ehrlich_et_al_2025; chen2021evaluating), the test repair LLM call uses a higher temperature (0.8) to encourage the model to try diverse solutions.
 We refer to the test generated at this phase as e-Otter test.
-(a) Issue description
-(b) Test patch
-Figure 3. Issue description and test patch for sympy__sympy-23413. Some lines are omitted from the patch due to space constraints.
 Issue Description Morphs and Test-Generation Context Masks
-Recent works both on generating SWE-patches (Ehrlich et al., 2025; Xia et al., 2025; Jain et al., 2025) and reproduction tests (Ahmed et al., 2025; Wang et al., 2024) have adopted inference scaling.
+Recent works both on generating SWE-patches (ehrlich_et_al_2025; xia_et_al_2025; jain_et_al_2025) and reproduction tests (ahmed_et_al_2025; wang_et_al_2024) have adopted inference scaling.
 Most works rely on higher temperatures to increase the diversity of solutions.
@@
 Also, in a chain of multiple LLM calls, using a higher temperature in each call exponentially increases the number of samples.
-Instead of higher temperature, Otter++ (Ahmed et al., 2025) attempted heterogeneous prompting via five masks hiding different parts of contexts and achieved a higher fail-to-pass rate @5.
+Instead of higher temperature, Otter++ (ahmed_et_al_2025) attempted heterogeneous prompting via five masks hiding different parts of contexts and achieved a higher fail-to-pass rate @5.
 We refer to the five tests from Otter++ masks as planner, full, testLoc, patchLoc, and none.
@@
 Note that in this augmentation process, the model may contaminate the issue from a memorized test.
-We manually validate the possibility of contamination by the LLM and found that in 2% of cases, we observed some possible contamination where the model might have used memorized data to resolve the problem (see details in Section 5.1).
+We manually validate the possibility of contamination by the LLM and found that in 2% of cases, we observed some possible contamination where the model might have used memorized data to resolve the problem (see details in Section 5).
 Also, we are not performing anything better than zero-shot prompting here, and we observed that zero-shot prompting is not as effective as other approaches.
@@
 The model may simply rewrite the issue with minor modifications.
+The motivation behind morphing is to change the issue such as to give the model a different perspective.
+Of course, this only happens if the human-written issue does not already have the characteristics that the morph would add.
+This raises the question how often that is the case; to answer it, we manually labeled human-written issues with morph categories.
+Upon reviewing the entire dataset (449 issues), only 28% are written in a standard format, which means our standardization morphing makes substantial changes to the remaining 72%.
+For other morphing techniques, substantial changes happen to 28% (simple), 71% (dropCode), 47% (initTest), and 90% (initCode) issues.
+There is certainly an opportunity to add even more diversity, but Figure 5 shows diminishing returns.
 We rerun the Otter pipeline using these new morphed issue descriptions and generate five new sets of tests.
@@
 Prompts and examples for execution-augmented test repair and issue description morphs are in the supplementary material.
+(a) Issue description
+(b) Test patch
+Figure 3. Issue description and test patch for sympy__sympy-23413. Some lines are omitted from the patch due to space constraints.
 3.2. Code Patch Generator
@@
 Prior work does something similar but in the other direction, using generated tests to help with code selection.
-CodeT (Chen et al., 2023) generates both code samples and corresponding test cases, then executes the generated code samples using the generated test cases and performs a dual execution agreement to choose the best code sample.
-Agentless (Xia et al., 2025) and later CodeMonkeys (Ehrlich et al., 2025) also use reproduction tests to select patches from multiple candidates but do not evaluate the quality of the reproduction test.
+CodeT (chen_et_al_2023) generates both code samples and corresponding test cases, then executes the generated code samples using the generated test cases and performs a dual execution agreement to choose the best code sample.
+Agentless (xia_et_al_2025) and later CodeMonkeys (ehrlich_et_al_2025) also use reproduction tests to select patches from multiple candidates but do not evaluate the quality of the reproduction test.
 Our e-Otter++ uses code patches to evaluate the quality of generated reproduction test candidates and attempt to choose the best one.
@@
   • RQ5: How does our approach help in validating code patches?
-We evaluate our approach on two benchmarks: TDD-Bench Verified [^2] https://github.com/IBM/TDD-Bench-Verified and SWT-bench Lite [^3] https://github.com/logic-star-ai/swt-bench.
-Both benchmarks are derived from the popular issue resolution benchmark, SWE-bench (Jimenez et al., 2024).
-We report the fail-to-pass ($F\!\!\to\!\!P$) rate to show the effectiveness of the approach following prior works (Mündler et al., 2024; Nashid et al., 2025; Ahmed et al., 2025).
+  • RQ6: How effective is e-Otter++ with a contamination-free benchmark?
+We primarily evaluate our approach on two benchmarks: TDD-Bench Verified [^1] https://github.com/IBM/TDD-Bench-Verified and SWT-bench Lite [^2] https://github.com/logic-star-ai/swt-bench.
+Both benchmarks are derived from the popular issue resolution benchmark, SWE-bench (jimenez_et_al_2024).
+We report the fail-to-pass ($F\!\!\to\!\!P$) rate to show the effectiveness of the approach following prior works (mundler_et_al_2024; nashid_et_al_2025; ahmed_et_al_2025).
 We have primarily used two models, GPT-4o and Claude-3.7-Sonnet.
-To have a fair comparison with Otter and Otter++, we used the same version of the GPT-4o model (gpt-4o-2024-08-06) as in their paper (Ahmed et al., 2025).
+To have a fair comparison with Otter and Otter++, we used the same version of the GPT-4o model (gpt-4o-2024-08-06) as in their paper (ahmed_et_al_2025).
+We also report results on SWE-rebench (badertdinov2025swe) with Claude-3.7-Sonnet to demonstrate its effectiveness on a contamination-free benchmark.
 4.1. Performance of e-Otter and e-Otter++ (RQ1)
@@
 For TDD-Bench Verified and SWT-bench Lite, the best baselines report 38.8% and 37.7%, respectively.
-Our Claude-3.7-Sonnet based e-Otter++ achieves state-of-the-art performance on both benchmarks, achieving 63.0% and 52.5% on TDD-Bench Verified and SWT-bench Lite, respectively.
-We perform McNemar’s test (McNemar, 1947) between our best test generator and all prior approaches and it significantly outperforms all the baselines on both TDD-Bench Verified and SWT-bench Lite with a 99% confidence interval ($p<0.01$).
+Our Claude-3.7-Sonnet based e-Otter++ achieves state-of-the-art performance on both benchmarks, achieving 63.0% and 52.5% on TDD-Bench Verified and SWT-bench Lite, respectively using CodePatch + Coverage as selector (see Table 9).
+We perform McNemar’s test (mcnemar1947note) between our best test generator and all prior approaches and it significantly outperforms all the baselines on both TDD-Bench Verified and SWT-bench Lite with a 99% confidence interval ($p<0.01$).
 McNemar’s test compares the performance of two systems on the same set of binary-labeled instances, highlighting statistically significant differences in their error patterns.
@@
 Although not directly comparable to AEGIS, e-Otter++ with Claude-3.7-Sonnet outperforms AEGIS on SWT-bench Lite with a 52.5% $F\!\!\to\!\!P$ rate.
-Note that, as mentioned in Section 2, the two benchmarks have different implementations.
-We primarily used the TDD-Bench harness because of its speed.
-We also ran the SWT-Bench harness and observed a slight drop in performance (50.7%) due to dependency issues and regression tests.
-While the SWT-bench paper (Mündler et al., 2024) only describes the Lite subset of their benchmark, the leaderboard recently also added a Verified subset.
-SWT-bench Verified is similar to TDD-Bench Verified, as both are derived from SWE-Bench-Verified (Chowdhury et al., 2024).
+While the SWT-bench paper (mundler_et_al_2024) only describes the Lite subset of their benchmark, the leaderboard recently also added a Verified subset.
+SWT-bench Verified is similar to TDD-Bench Verified, as both are derived from SWE-Bench-Verified (chowdhury_et_al_2024).
 Although both benchmarks have some non-overlapping samples, systems usually perform similarly on them.
@@
 Table 5 shows that GPT-4o achieves a 0.78 to 0.83 F-score as a critic, which indicates the benefit of using LLMs as critics.
+Golden tests have no false positives, resulting in a precision of 1.
+However, they can have false negatives (if the model mistakenly indicates that the test is failing for the wrong reason).
 Figure 4. Number of attempts taken in test repair
@@
 Morphs and masks yield heterogeneous prompts, which in turn increase the diversity of candidate tests, thus hopefully increasing $F\!\!\to\!\!P$ @ N.
-We reuse 5 heterogeneous prompts including masks from Otter++ (Ahmed et al., 2025) and add 5 more prompts using morphs for a total of 10 heterogeneous prompts.
+We reuse 5 heterogeneous prompts including masks from Otter++ (ahmed_et_al_2025) and add 5 more prompts using morphs for a total of 10 heterogeneous prompts.
 Table 6 shows that morphs increase $F\!\!\to\!\!P$ @ N, even without masks, both before (Otter) and after (e-Otter) test repair.
@@
 Why not use high temperature instead of heterogeneous prompting?
-Prior works (Xia et al., 2025; Ehrlich et al., 2025; Chen et al., 2021) mostly depend on high temperature for multi-sampling.
+Prior works (xia_et_al_2025; ehrlich_et_al_2025; chen2021evaluating) mostly depend on high temperature for multi-sampling.
 Figure 5 compares the results of heterogeneous prompting with the samples generated with high temperature.
-We tried three distinct temperatures (0.5, 0.8, and 1.0) and find that the best results are obtained at a temperature of 0.8, which is consistent with prior research (Chen et al., 2021).
+We tried three distinct temperatures (0.5, 0.8, and 1.0) and find that the best results are obtained at a temperature of 0.8, which is consistent with prior research (chen2021evaluating).
 However, our approach, “plan + mask + morph”, significantly outperforms samples generated with high temperature.
@@
 This indicates that our approaches are also applicable to open-source models.
+Table 8 shows how much a morph can increase $F\!\!\to\!\!P$ @ 2 from $F\!\!\to\!\!P$ @ 1 (using base e-Otter) for TDD-bench-Verified because multiple morphs can independently benefit a single issue.
+Each morph increases $F\!\!\to\!\!P$ by 29%-32% for Claude and 11%-21% for GPT-4o.
+Table 8. Contribution of each prompt to $F\!\!\to\!\!P$ @ N
+Variants Claude-3.7-Sonnet GPT-4o
+$F\!\!\to\!\!P$ @ 2 Change $F\!\!\to\!\!P$ @ 2 Change
+Standard 235 +31.28% 194 +16.16%
+Simple 235 +31.28% 186 +11.37%
+dropCode 231 +29.05% 199 +19.16%
+initTest 228 +27.37% 197 +17.96%
+initPatch 236 +31.84% 203 +21.55%
 Finding 3. The samples with “plan + mask + morph” increase $F\!\!\to\!\!P$ @ N by more than 13% compared to “plan + mask”, regardless of the model and benchmark.
@@
 Ahmed et al.
-(Ahmed et al., 2025) took the top 22 systems from the leaderboard and ran the five tests generated by Otter++.
+(ahmed_et_al_2025) took the top 22 systems from the leaderboard and ran the five tests generated by Otter++.
 They filtered out a code patch if all the tests failed on it.
@@
 Mündler et al.
-(Mündler et al., 2024) performed a similar experiment with one system.
+(mundler_et_al_2024) performed a similar experiment with one system.
 Note that neither of these two prior approaches could achieve a $F\!\!\to\!\!P$ rate beyond 50% even with 5 samples.
@@
 Finding 5. We achieve 47.9% to 72.8% recall, regardless of the model, while increasing the precision of the SWE-patches.
+4.6. Performance on SWE-rebench (RQ6)
+To show that our proposed approach is effective for non-contaminated models, we extend our evaluation on SWE-rebench with the Claude-3.7-Sonnet model.
+The training data cut-off date for Claude-3.7-Sonnet is November 2024 (claude-cutoff).
+To address data contamination, we consider instances from SWE-rebench (badertdinov2025swe) with a creation date in 2025 and check whether the developers’ written golden tests go from fail-to-pass to select the instances.
+This experiment uses 267 instances.
+Table 11 shows that execution-augmented test repair increases the fail-to-pass rate from 21% to 72% for all prompt variants.
+Table 12 also shows that heterogeneous prompting increases the fail-to-pass @ N by more than 32%, which is consistent with our findings on other benchmarks.
+The last step of our pipeline, test selection, uses surrogate patches.
+We use leaderboard patches shared by the SWE-rebench authors instead of Agentless to evaluate this step.
+The average resolve rate of these patches is less than 40%, and these patches are also free from contamination.
+Table 13 shows that e-Otter++ achieved a 42.7% $F\!\!\to\!\!P$ rate on SWE-rebench, which is 75% higher than Otter.
+The absolute numbers on the non-contaminated SWE-rebench are notably lower than on TDD-Bench or SWT-bench, and contamination might be a major factor behind that.
+Another factor is that the samples from SWE-rebench come from 152 projects, most of which are less popular as the projects in SWE-bench.
+Popular projects are relatively older, larger, and more likely to be included in training data.
+Despite lower absolute numbers, our approach shows a great relative improvement, demonstrating that our two primary contributions also work for contamination-free benchmarks like SWE-rebench.
+Table 11. Effectiveness of execution feedback on SWE-rebench
+Prompt Otter e-Otter Change
+# of F-P in % # of F-P in %
+Planner 65 24.3 86 32.2 32.3%
+Context Masking full 56 21.0 71 26.6 26.8%
+testLoc 49 18.4 67 25.1 36.7%
+patchLoc 41 15.4 60 22.5 46.3%
+none 25 9.4 43 16.1 72.0%
+Issue Morphing standard 60 22.5 73 27.3 21.7%
+simple 45 16.9 68 25.5 51.1%
+dropcode 49 18.4 75 28.1 53.1%
+initTests 50 18.7 72 27.0 44.0%
+initPatch 56 21.0 70 26.2 25.0%
+Table 12. Impact of heterogeneous prompts on SWE-rebench
+Category N Otter e-Otter
+$F\!\!\to\!\!P$ @ N In % Change $F\!\!\to\!\!P$ @ N In % Change
+plan + mask 5 97 36.3 NA 119 44.6 NA
+plan + morph 6 115 43.1 18.6% 146 54.7 22.7%
+plan + mask + morph 10 129 48.3 33.0% 158 59.2 32.8%
+Table 13. Performance of e-Otter++ on SWE-rebench
+Approach # of $F\!\!\to\!\!P$ in % Change from Otter
+Otter 65 24.3 NA
+Otter++ 87 32.6 33.8%
+e-Otter 86 32.2 32.3%
+e-Otter++ 114 42.7 75.4%
+Finding 6. e-Otter++ is effective regardless of the data contamination issue. It improves the $F\!\!\to\!\!P$ rate by a relative 75% compared to Otter on SWE-rebench.
 5. Discussion
 This section briefly discusses model contamination, test coverage, and a detailed analysis of the tests.
-5.1. Contamination through Issue Morphs
+Contamination through Issue Morphs.
 Test generation from issues could potentially suffer from model memorization and contamination since the model might have seen the data during the training phase.
 Both prior works by Ahmed et al.
-(Ahmed et al., 2025) and Mündler et al.
-(Mündler et al., 2024) addressed this issue and show that the model-generated tests are very different from developer-written tests and the impact of contamination is minimal.
-Since we build upon Ahmed et al.’s Otter++, we refer to their contamination analysis except for places where e-Otter++ might potentially introduce new contamination.
+(ahmed_et_al_2025) and Mündler et al.
+(mundler_et_al_2024) addressed this issue and show that the model-generated tests are very different from developer-written tests and the impact of contamination is minimal.
+However, e-Otter++ might potentially introduce new contamination.
 Specifically, issue morphs might introduce contamination while rewriting the issue.
@@
 Table 14 shows the inter-rater agreement and outcome of the study.
-We use Fleiss’ Kappa (Fleiss, 1971) to compute inter-rater agreement, and the raters show fair to substantial agreement [^4] Kappa ¡ 0: Poor, 0.01–0.20: Slight, 0.21–0.40: Fair, 0.41–0.60: Moderate, 0.61–0.80: Substantial, 0.81–1.00: Almost perfect agreement..
+We use Fleiss’ Kappa (fleiss1971measuring) to compute inter-rater agreement, and the raters show fair to substantial agreement [^3] Kappa ¡ 0: Poor, 0.01–0.20: Slight, 0.21–0.40: Fair, 0.41–0.60: Moderate, 0.61–0.80: Substantial, 0.81–1.00: Almost perfect agreement..
 Note that for contamination, we observe a very low kappa value because all raters mostly choose the same category.
-This phenomenon is known as the kappa paradox (Feinstein and Cicchetti, 1990).
+This phenomenon is known as the kappa paradox (feinstein1990high).
 Overall, 40% of modified issues have some information addition and 68% of issues have some information removal.
@@
 But whether this case constitutes actual contamination is unclear since that line is present in the original issue description.
-5.2. Coverage of Generated Tests
+Coverage of the Generated Tests.
 Coverage is an important metric to measure the adequacy of tests.
@@
 Figure 7 shows that $F\!\!\to\!\!P$ tests have higher coverage compared to other tests.
-These findings are consistent with prior works (Ahmed et al., 2025; Mündler et al., 2024) and motivate using coverage as a tie-breaker for test selection (of course, test selection uses coverage on generated code patches, not golden ones).
+These findings are consistent with prior works (ahmed_et_al_2025; mundler_et_al_2024) and motivate using coverage as a tie-breaker for test selection (of course, test selection uses coverage on generated code patches, not golden ones).
 Figure 7. Coverage of the generated tests
-5.3. Detailed Analysis of the Tests
+Detailed Analysis of the Tests.
 Table15 presents an analysis of the e-Otter-generated tests using GPT-4o from different perspectives.
@@
 However, we found 43 samples in e-Otter where the test passed on $c_{\textrm{old}}$.
-Also, our analysis shows that tests with assertion failures have a higher success rate (50.4% fail-to-pass rate) compared to other groups.
+Also, our analysis shows that tests with assertion failures have a higher success rate (50.4% $F\!\!\to\!\!P$ rate) compared to other groups.
 Table 15. Analysis of the e-Otter-generated tests using GPT-4o model from different perspectives.
@@
 Error 88 8 9.1
-6. Threats to Validity
-One limitation of our work is that it only experiments with Python.
-As a result, the findings may not generalize to other programming languages or repositories outside the benchmarks.
-However, it is worth noting that prior works (Jimenez et al., 2024; Xia et al., 2025; Zhang et al., 2024), which share this same limitation, have still been impactful and contributed significantly to the field.
-One of the key concerns with SWE-bench related work is the contamination or memorization problem of the model.
-The model may have seen the data during the training process, and that might positively impact the results.
-Prior works (Jimenez et al., 2024; Ahmed et al., 2025; Mündler et al., 2024) have attempted to address this problem.
-The SWE-bench paper claims that “difficulty does not correlate with issue resolution date”, indicating minimal contamination.
-Ahmed et al.
-show that Otter-generated tests differ substantially from developer-generated tests.
-In this work, we reuse Otter-generated tests.
-The only way we can introduce new contamination is while rewriting the issue description for morphs.
-However, Section 5.1 demonstrates through manual validation that issue morphs do not introduce contamination.
-We believe our results are not influenced by major model contamination.
+Cost and Time.
 Our system, e-Otter++, uses both sequential inference scaling (via test repair) and parallel inference scaling (via heterogeneous prompting).
-This increases the number of LLM calls.
 Furthermore, we also use Agentless code patches for test selection, which requires several LLM calls.
-We did not generate the code patches for Agentless ourselves; instead, we relied on code patches shared by the authors for both the Claude and GPT-4o models.
+We did not generate the code patches for Agentless ourselves; instead, we relied on code patches shared by the authors for both models.
 However, based on their reported numbers and our experiments, we can estimate the cost for each sample.
@@
 In that case, the additional cost of generating code patches will be amortized.
+Additionally, excluding Agentless, e-Otter++ costs \$1.70 (\$0.17/candidate) for Claude and \$1.10 (\$0.11/candidate) for GPT-4o.
+Figure 5 shows how the $F\!\!\to\!\!P$ @ N changes with the number of candidates.
+For example, 6 samples yield 230 $F\!\!\to\!\!P$ tests and generating 6 samples costs \$0.66.
+It is hard to report the end-to-end time because several steps of our process can be parallelized.
+We did most of our experiments sequentially due to shared resources.
+Docker setup takes 5-10 minutes based on the repository.
+Other steps (i.e., test variant generation, agentless generation, ranker) can take up to 2–3 minutes, some of which are parallelizable.
+6. Threats to Validity
+One limitation of our work is that it only experiments with Python.
+As a result, the findings may not generalize to other programming languages.
+However, it is worth noting that prior works (jimenez_et_al_2024; xia_et_al_2025; zhang_et_al_2024), which share this same limitation, have still been impactful and contributed significantly to the field.
+One of the key concerns with SWE-bench related work is the contamination or memorization problem of the model.
+To address this problem, we evaluated e-Otter++ on SWE-rebench and observed similar impact from execution-augmented test repair and heterogeneous prompting.
 Another limitation of e-Otter++ is that it considers only one test file and generates only one block of code.
@@
 Despite that, e-Otter++ exceeds the state-of-the-art performance, so we leave further improvements to future work.
+It is hard to ensure the correctness of the generated tests.
+TDD-bench-Verified and SWT-bench use coverage in conjunction with the $F\!\!\to\!\!P$ property to be considered a success.
+Coverage is a primary metric for test quality.
+These benchmarks allow us to apply the golden patch (written by real developers and likely to be correct), run the test, and check both how well it covers the edited lines and whether it goes from fail to pass.
+While this is no correctness guarantee, it is a strong proxy for correctness and followed by prior works (mundler_et_al_2024; ahmed_et_al_2025).
+Besides, using imperfect patches or tests comes with certain risks, but prior work (chen_et_al_2023) has shown that dual execution agreement can help discard samples with potential errors.
+To further mitigate risk, we can use separate systems for surrogate patches and the proposed final patch (of course, with increased cost).
 7. Related Work
 There has been a flurry of recent papers about generating reproduction tests from issues, which is the same problem we tackle in our paper.
-Libro generates tests for Java bug reports (Kang et al., 2023).
+Libro generates tests for Java bug reports (kang_yoon_yoo_2023).
 It repairs tests based on compiler feedback, not execution feedback like we do, and generates several tests using high LLM decoding temperature, not morphs and masks like we do.
 Plein et al.
-also explore the feasibility of generating tests for Java bug reports, albeit without test repair (Plein et al., 2024).
+also explore the feasibility of generating tests for Java bug reports, albeit without test repair (plein_et_al_2024).
 Mündler et al.
-(Mündler et al., 2024) introduce SWE-Agent+, a Python reproduction test generation agent derived from the SWE-Agent (Yang et al., 2024) code patch generator.
+(mundler_et_al_2024) introduce SWE-Agent+, a Python reproduction test generation agent derived from the SWE-Agent (yang_et_al_2024) code patch generator.
 While SWE-Agent+ has access to tools that can give it execution feedback, it leverages such feedback less effectively than e-Otter++.
-The same holds for EvoCoder (Lin et al., 2024), also derived from SWE-Agent, and for USEAgent (Applis et al., 2025), derived from AutoCodeRover (Zhang et al., 2024).
-AEGIS is a novel multi-agent system for reproduction test generation (Wang et al., 2024).
+The same holds for EvoCoder (lin_et_al_2024), also derived from SWE-Agent, and for USEAgent (applis_et_al_2025), derived from AutoCodeRover (zhang_et_al_2024).
+AEGIS is a novel multi-agent system for reproduction test generation (wang_et_al_2024).
 It repairs tests based on execution feedback from $c_{\textrm{old}}$, but unlike our work, does not augment the test repair prompt.
 Also, it standardizes issue descriptions, but unlike our work, does not use morphs or masks to increase test candidate pool diversity.
-Otter++ (Ahmed et al., 2025) is an LLM-based workflow for reproduction test generation that we use as a component in e-Otter++, which extends it in three important ways: (i) augmenting the test repair prompt; (ii) using morphs (Otter++ only uses masks); and (iii) leveraging execution feedback from generated candidate code patches.
+Otter++ (ahmed_et_al_2025) is an LLM-based workflow for reproduction test generation that we use as a component in e-Otter++, which extends it in three important ways: (i) augmenting the test repair prompt; (ii) using morphs (Otter++ only uses masks); and (iii) leveraging execution feedback from generated candidate code patches.
 Our experiments show that these novel contributions greatly improve the $F\!\!\to\!\!P$ rate.
-BRT Agent is a new agent for generating reproduction tests (Cheng et al., 2025), and Issue2Test is a new non-agentic workflow for the same task (Nashid et al., 2025); neither augments test repair prompts, uses masks or morphs, or leverages candidate code patches.
+BRT Agent is a new agent for generating reproduction tests (cheng_et_al_2025), and Issue2Test is a new non-agentic workflow for the same task (nashid_et_al_2025); neither augments test repair prompts, uses masks or morphs, or leverages candidate code patches.
 Unlike e-Otter++ and the above papers on using generated code patches to help pick or improve reproduction tests, there is also work on the opposite: using reproduction tests to help pick or improve code patches.
-Papers on using tests to help pick (but not improve) code patches include Smith et al.’s 2015 paper (Smith et al., 2015), CodeT (Chen et al., 2023), Agentless (Xia et al., 2025), and R2E-Gym (Jain et al., 2025).
-Similarly, several papers use tests to help improve (but not pick) code patches (Zhang et al., 2024; Yang et al., 2024; Chen et al., 2024; Arora et al., 2024; Ruan et al., 2024; Mathews and Nagappan, 2024).
-Some papers even do both, using tests to help pick and improve code patches: CodeMonkeys (Ehrlich et al., 2025) runs several parallel workflows, each of which iteratively refines a $\langle c_{\mathrm{new}},y\rangle$ pair; and PatchPilot (Li et al., 2025) uses reproduction tests for bug localization before code patch generation, and later uses them again to pick among multiple candidate code patches.
+Papers on using tests to help pick (but not improve) code patches include Smith et al.’s 2015 paper (smith_et_al_2015), CodeT (chen_et_al_2023), Agentless (xia_et_al_2025), and R2E-Gym (jain_et_al_2025).
+Similarly, several papers use tests to help improve (but not pick) code patches (zhang_et_al_2024; yang_et_al_2024; chen_et_al_2024; arora_et_al_2024; ruan_zhang_roychoudhury_2024; mathews_nagappan_2024).
+Some papers even do both, using tests to help pick and improve code patches: CodeMonkeys (ehrlich_et_al_2025) runs several parallel workflows, each of which iteratively refines a $\langle c_{\mathrm{new}},y\rangle$ pair; and PatchPilot (li_et_al_2025) uses reproduction tests for bug localization before code patch generation, and later uses them again to pick among multiple candidate code patches.
 However, while most of the work described in this paragraph uses generated tests, test generation is not their main focus: their reproduction test generation components are more basic, and they lack empirical evaluation of the $F\!\!\to\!\!P$ rate of generated tests.
@@
 All our generated tests and logs are available at https://doi.org/10.5281/zenodo.15860535.
+Acknowledgements.
+We gratefully acknowledge Ibragim Badertdinov and Chunqiu Steven Xia for sharing their model-generated SWE-rebench and Agentless patches.

### v3

@@
 Unfortunately, most issues do not come with functioning reproduction tests, so this paper explores how to generate them automatically.
-The primary challenge in this setting is that the code to be tested is either missing or wrong, as evidenced by the existence of the issue in the first place.
-This has held back test generation for this setting: without the correct code to execute, it is difficult to leverage execution feedback to generate good tests.
-This paper introduces novel techniques for leveraging execution feedback to get around this problem, implemented in a new reproduction test generator called e-Otter++.
+One technique that helps with this task is inference scaling, but traditional temperature-based scaling tends to lack diversity.
+This paper introduces heterogeneous prompting to address this problem.
+Another technique that helps with this task is execution feedback, but this is hampered by the fact that the new code (after issue resolution) to execute does not yet exist.
+This paper introduces novel approaches to get around that problem.
+We implemented our techniques in a new reproduction test generator called e-Otter++.
 Experiments show that e-Otter++ represents a leap ahead in the state-of-the-art for this problem, generating tests with an average fail-to-pass rate of 63% on the TDD-Bench Verified benchmark.
 LLMs, SWE Patches, Reproduction Tests
-journalyear: 2026copyright: rightsretainedconference: 2026 IEEE/ACM 46th International Conference on Software Engineering; April 15–17, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06
+journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3787837isbn: 979-8-4007-2025-3/2026/04
 1. Introduction
@@
 But such tests, when they exist, are invaluable for software development.
-Before issue resolution, tests can reproduce the issue to confirm whether it is real, make it more concrete, and help stakeholders agree upon the desired behavior—following the highly regarded software development practice of test-driven development (TDD) (beck_2002).
+Before issue resolution, tests can reproduce the issue to confirm whether it is real, make it more concrete, and help stakeholders agree upon the desired behavior—following the highly regarded software development practice of test-driven development (TDD) (Beck, 2002).
 Once a code change (a patch) is proposed to resolve the issue, running the same reproduction tests helps decide whether to accept the patch.
@@
 Unfortunately, generating $F\!\!\to\!\!P$ tests is challenging, because issue descriptions are informal and because the new code patch on which the tests should pass does not yet exist.
-In addition to helping human software engineers, generating tests from SWE issues can also assist automated SWE agents, as several SWE agents also rely on reproduction tests (arora_et_al_2024; chen_et_al_2024; ehrlich_et_al_2025; jain_et_al_2025; li_et_al_2025; ruan_zhang_roychoudhury_2024; xia_et_al_2025; yang_et_al_2024; zhang_et_al_2024).
+In addition to helping human software engineers, generating tests from SWE issues can also assist automated SWE agents, as several SWE agents also rely on reproduction tests (Arora et al., 2024; Chen et al., 2024; Ehrlich et al., 2025; Jain et al., 2025b; Li et al., 2025; Ruan et al., 2024; Xia et al., 2025; Yang et al., 2024; Zhang et al., 2024).
 There have been several recent attempts at using large language models (LLMs) to automatically generate $F\!\!\to\!\!P$ tests from SWE issues, without access to the resolving code patches.
-Libro uses an LLM to directly generate 50 tests from Java bug reports, then performs rule-based postprocessing and ranking (kang_yoon_yoo_2023).
-Other attempts focus on Python, modifying a SWE agent designed to automatically generate code patches to generate tests instead (lin_et_al_2024; mundler_et_al_2024).
-Recent papers proposed novel LLM-based flows or agents custom-tailored to the problem of generating tests from issues, including AEGIS (wang_et_al_2024), Otter++ (ahmed_et_al_2025), BRT agent (cheng_et_al_2025), and Issue2Test (nashid_et_al_2025).
+Libro uses an LLM to directly generate 50 tests from Java bug reports, then performs rule-based postprocessing and ranking (Kang et al., 2023).
+Other attempts focus on Python, modifying a SWE agent designed to automatically generate code patches to generate tests instead (Lin et al., 2024; Mündler et al., 2024).
+Recent papers proposed novel LLM-based flows or agents custom-tailored to the problem of generating tests from issues, including AEGIS (Wang et al., 2024), Otter++ (Ahmed et al., 2025), BRT agent (Cheng et al., 2025), AssertFlip (Khatib et al., 2025), and Issue2Test (Nashid et al., 2025).
 Table 1 shows how e-Otter++ is procedurally distinct.
 Note that Otter, Issue2Test, and e-Otter do not use inference scaling or test selection.
-There are two benchmarks for generating tests from issues: SWT-bench (mundler_et_al_2024) (derived from SWE-bench (jimenez_et_al_2024)) and TDD-Bench Verified (ahmed_et_al_2024-tdd) (derived from SWE-bench Verified (chowdhury_et_al_2024)).
+There are two benchmarks for generating tests from issues: SWT-bench (Mündler et al., 2024) (derived from SWE-bench (Jimenez et al., 2024)) and TDD-Bench Verified (Ahmed et al., 2024) (derived from SWE-bench Verified (Chowdhury et al., 2024)).
 SWT-bench Verified, also derived from SWE-bench Verified, is very similar to TDD-Bench Verified.
-SWT-bench has a leaderboard (https://swtbench.com), topped by the unpublished Amazon Q Developer Agent with an $F\!\!\to\!\!P$ rate of 37.7% on SWT-bench Lite and 49.0% on SWT-bench Verified.
+SWT-bench has a leaderboard (https://swtbench.com), topped (as of when we submitted this paper) by the unpublished Amazon Q Developer Agent with an $F\!\!\to\!\!P$ rate of 37.7% on SWT-bench Lite and 49.0% on SWT-bench Verified.
 However, as there is no paper about Amazon Q Developer Agent, its inner workings are not publicly known.
-The top system on SWT-bench Verified with a published paper is Otter++ (ahmed_et_al_2025) with an $F\!\!\to\!\!P$ rate of 37.0%, and it also has the exact same $F\!\!\to\!\!P$ rate of 37.0% on TDD-Bench Verified.
+The top system on SWT-bench Verified with a published paper is Otter++ (Ahmed et al., 2025) with an $F\!\!\to\!\!P$ rate of 37.0%, and it also has the exact same $F\!\!\to\!\!P$ rate of 37.0% on TDD-Bench Verified.
 Given how similar the two verified benchmarks are, this paper evaluates on TDD-Bench Verified only.
@@
 The “e-” refers to execution feedback: e-Otter++ outperforms prior work by leveraging execution feedback more effectively.
-Additionally, we apply our approach to the contamination-free benchmark SWE-rebench (badertdinov2025swe) to show the effectiveness of our two primary contributions (execution feedback and heterogeneous prompting), as SWE-bench is suspected to be exposed to recent frontier models during training (liang2025swe).
+The other core innovation in e-Otter++ is heterogeneous prompting.
+Additionally, we apply our approach to the contamination-free benchmark SWE-rebench (Badertdinov et al., 2025) to show the generalizability of our innovations, as SWE-bench is suspected to be exposed to recent frontier models during training (Liang et al., 2025a).
 The first place where e-Otter++ uses execution feedback is after generating a candidate test.
@@
 Ideally, this test pool should be diverse, and most prior work turns up the LLM decoding temperatures to increase candidate diversity.
-Unlike prior work, e-Otter++ instead gets diversity from two kinds of systematic prompt variations: issue description morphs, which use an LLM to rewrite the issue text in various ways before test generation, and prompt context masks, which elide various subsets of features from the test-generation prompt.
+Unlike prior work, e-Otter++ instead gets diversity from heterogeneous prompting.
+It uses two kinds of systematic prompt variations: issue description morphs, which use an LLM to rewrite the issue text in various ways before test generation, and prompt context masks, which elide various subsets of features from the test-generation prompt.
 Our experiments show that morphs and masks greatly increase the likelihood of at least one candidate test being $F\!\!\to\!\!P$.
@@
 2. Problem Statement and Background
-This section discusses the problem statement and briefly introduces prior work we build upon: Otter (ahmed_et_al_2025), Agentless (xia_et_al_2025), SWT-bench (mundler_et_al_2024), and TDD-Bench Verified (ahmed_et_al_2024-tdd).
+This section discusses the problem statement and briefly introduces prior work we build upon: Otter (Ahmed et al., 2025), Agentless (Xia et al., 2025), SWT-bench (Mündler et al., 2024), and TDD-Bench Verified (Ahmed et al., 2024).
 Figure 1.
@@
 2.2. Reproduction Test Generation: Otter
-This paper introduces e-Otter++, which extends an earlier system called Otter++ (ahmed_et_al_2025).
+This paper introduces e-Otter++, which extends an earlier system called Otter++ (Ahmed et al., 2025).
 Both e-Otter++ and Otter++ generate tests directly from issue descriptions and $c_{\textrm{old}}$.
@@
 One of the new techniques in e-Otter++ is how it uses execution feedback from automatically generated candidate code patches.
-Rather than implementing a new code patch generator, it reuses an existing one, Agentless (xia_et_al_2025), as a component.
+Rather than implementing a new code patch generator, it reuses an existing one, Agentless (Xia et al., 2025), as a component.
 Agentless implements a hierarchical bug localization process.
@@
 Our new system e-Otter++ uses only the first part of Agentless, without the test generation or patch selection of Agentless, as a component (see Section 3.2).
-While we could have adopted any code patch generator, we chose Agentless because it is open-source and candidate patches are publicly available (agentless_artifact).
+While we could have adopted any code patch generator, we chose Agentless because it is open-source and candidate patches are publicly available (1).
 2.4. Test Evaluation Benchmarks
 There are two benchmarks for evaluating bug reproduction tests.
-SWT-bench Lite (mundler_et_al_2024) was derived primarily from the SWE-bench Lite dataset (jimenez_et_al_2024) and contains 276 samples.
-TDD-Bench Verified (ahmed_et_al_2024-tdd) was derived from SWE-bench Verified (chowdhury_et_al_2024).
+SWT-bench Lite (Mündler et al., 2024) was derived primarily from the SWE-bench Lite dataset (Jimenez et al., 2024) and contains 276 samples.
+TDD-Bench Verified (Ahmed et al., 2024) was derived from SWE-bench Verified (Chowdhury et al., 2024).
 In TDD-Bench Verified, the authors ran 500 samples from SWE-bench Verified, filtered out samples with zero test coverage or those violating the $F\!\!\to\!\!P$ property of the tests, and ended up with 449 samples.
@@
 Chen et al.
-(chen_ma_jiang_2025) reported that Python execution errors during the issue resolution phase correlate with lower resolution rates.
+(Chen et al., 2025b) reported that Python execution errors during the issue resolution phase correlate with lower resolution rates.
 They also identified the most prevalent errors—such as ModuleNotFoundError and TypeError—in the generated code, which tend to differ greatly from the reason mentioned in the issue description.
@@
 Then, the repair loop starts over from this new test as $y$, until the critic is satisfied or for a maximum of 10 iterations.
-Following prior works (xia_et_al_2025; ehrlich_et_al_2025; chen2021evaluating), the test repair LLM call uses a higher temperature (0.8) to encourage the model to try diverse solutions.
+Following prior works (Xia et al., 2025; Ehrlich et al., 2025; Chen et al., 2021), the test repair LLM call uses a higher temperature (0.8) to encourage the model to try diverse solutions.
 We refer to the test generated at this phase as e-Otter test.
 Issue Description Morphs and Test-Generation Context Masks
-Recent works both on generating SWE-patches (ehrlich_et_al_2025; xia_et_al_2025; jain_et_al_2025) and reproduction tests (ahmed_et_al_2025; wang_et_al_2024) have adopted inference scaling.
+Recent works both on generating SWE-patches (Ehrlich et al., 2025; Xia et al., 2025; Jain et al., 2025b) and reproduction tests (Ahmed et al., 2025; Wang et al., 2024) have adopted inference scaling.
 Most works rely on higher temperatures to increase the diversity of solutions.
@@
 Also, in a chain of multiple LLM calls, using a higher temperature in each call exponentially increases the number of samples.
-Instead of higher temperature, Otter++ (ahmed_et_al_2025) attempted heterogeneous prompting via five masks hiding different parts of contexts and achieved a higher fail-to-pass rate @5.
+Instead of higher temperature, Otter++ (Ahmed et al., 2025) attempted heterogeneous prompting via five masks hiding different parts of contexts and achieved a higher fail-to-pass rate @5.
 We refer to the five tests from Otter++ masks as planner, full, testLoc, patchLoc, and none.
@@
 However, our goal is to increase the $F\!\!\to\!\!P$ rate @ N, which could be achieved by such rewriting.
-Prompts and examples for execution-augmented test repair and issue description morphs are in the supplementary material.
+Prompts and examples for execution-augmented test repair and issue description morphs are in our repository (2).
 (a) Issue description
@@
 Prior work does something similar but in the other direction, using generated tests to help with code selection.
-CodeT (chen_et_al_2023) generates both code samples and corresponding test cases, then executes the generated code samples using the generated test cases and performs a dual execution agreement to choose the best code sample.
-Agentless (xia_et_al_2025) and later CodeMonkeys (ehrlich_et_al_2025) also use reproduction tests to select patches from multiple candidates but do not evaluate the quality of the reproduction test.
+CodeT (Chen et al., 2023) generates both code samples and corresponding test cases, then executes the generated code samples using the generated test cases and performs a dual execution agreement to choose the best code sample.
+Agentless (Xia et al., 2025) and later CodeMonkeys (Ehrlich et al., 2025) also use reproduction tests to select patches from multiple candidates but do not evaluate the quality of the reproduction test.
 Our e-Otter++ uses code patches to evaluate the quality of generated reproduction test candidates and attempt to choose the best one.
@@
 We primarily evaluate our approach on two benchmarks: TDD-Bench Verified [^1] https://github.com/IBM/TDD-Bench-Verified and SWT-bench Lite [^2] https://github.com/logic-star-ai/swt-bench.
-Both benchmarks are derived from the popular issue resolution benchmark, SWE-bench (jimenez_et_al_2024).
-We report the fail-to-pass ($F\!\!\to\!\!P$) rate to show the effectiveness of the approach following prior works (mundler_et_al_2024; nashid_et_al_2025; ahmed_et_al_2025).
+Both benchmarks are derived from the popular issue resolution benchmark, SWE-bench (Jimenez et al., 2024).
+We report the fail-to-pass ($F\!\!\to\!\!P$) rate to show the effectiveness of the approach following prior works (Mündler et al., 2024; Nashid et al., 2025; Ahmed et al., 2025).
 We have primarily used two models, GPT-4o and Claude-3.7-Sonnet.
-To have a fair comparison with Otter and Otter++, we used the same version of the GPT-4o model (gpt-4o-2024-08-06) as in their paper (ahmed_et_al_2025).
-We also report results on SWE-rebench (badertdinov2025swe) with Claude-3.7-Sonnet to demonstrate its effectiveness on a contamination-free benchmark.
+To have a fair comparison with Otter and Otter++, we used the same version of the GPT-4o model (gpt-4o-2024-08-06) as in that paper (Ahmed et al., 2025).
+We also report results on SWE-rebench (Badertdinov et al., 2025) with Claude-3.7-Sonnet to demonstrate its effectiveness on a contamination-free benchmark.
 4.1. Performance of e-Otter and e-Otter++ (RQ1)
@@
 Benchmark Approach $F\!\!\to\!\!P$ In (%)
-TDD-Bench
-Verified (449
-samples)
-Zero-shot (GPT-4o) 84 18.7
-Otter++ (GPT 4o) 166 37.0
+TDD-Bench Verified (449 samples) Zero-shot (GPT-4o) 84 18.7
+Otter++ (GPT-4o) 166 37.0
 Otter++ (Claude-3.7-Sonnet) 174 38.8
@@
 e-Otter++ (Claude-3.7-Sonnet) 283 63.0
-SWT-bench
-Lite (276
-samples)
-ZeroShot (GPT-4) 16 5.8
+SWT-bench Lite (276 samples) ZeroShot (GPT-4) 16 5.8
 LIBRO (GPT-4) 42 15.2
@@
 SWE-Agent+ (GPT-4) 53 19.2
-Otter++ (GPT 4o) 80 29.0
+Otter++ (GPT-4o) 80 29.0
 Issue2Test (GPT-4o-mini) 84 30.4
 Otter++ (Claude-3.7-Sonnet) 84 30.4
+AssertFlip (GPT-4o) 99 36.0
 Amazon Q Developer Agent* (Unknown) 104 37.7
@@
 For TDD-Bench Verified and SWT-bench Lite, the best baselines report 38.8% and 37.7%, respectively.
-Our Claude-3.7-Sonnet based e-Otter++ achieves state-of-the-art performance on both benchmarks, achieving 63.0% and 52.5% on TDD-Bench Verified and SWT-bench Lite, respectively using CodePatch + Coverage as selector (see Table 9).
-We perform McNemar’s test (mcnemar1947note) between our best test generator and all prior approaches and it significantly outperforms all the baselines on both TDD-Bench Verified and SWT-bench Lite with a 99% confidence interval ($p<0.01$).
+Our Claude-3.7-Sonnet based e-Otter++ achieves state-of-the-art performance on both benchmarks, achieving 63.0% and 52.5% on TDD-Bench Verified and SWT-bench Lite, respectively, using CodePatch + Coverage as selector (see Table 8).
+We perform McNemar’s test (McNemar, 1947) between our best test generator and all prior approaches and it significantly outperforms all the baselines on both TDD-Bench Verified and SWT-bench Lite with a 99% confidence interval ($p<0.01$).
 McNemar’s test compares the performance of two systems on the same set of binary-labeled instances, highlighting statistically significant differences in their error patterns.
@@
 Although not directly comparable to AEGIS, e-Otter++ with Claude-3.7-Sonnet outperforms AEGIS on SWT-bench Lite with a 52.5% $F\!\!\to\!\!P$ rate.
-While the SWT-bench paper (mundler_et_al_2024) only describes the Lite subset of their benchmark, the leaderboard recently also added a Verified subset.
-SWT-bench Verified is similar to TDD-Bench Verified, as both are derived from SWE-Bench-Verified (chowdhury_et_al_2024).
+While the SWT-bench paper (Mündler et al., 2024) only describes the Lite subset of their benchmark, the leaderboard recently also added a Verified subset.
+SWT-bench Verified is similar to TDD-Bench Verified, as both are derived from SWE-Bench-Verified (Chowdhury et al., 2024).
 Although both benchmarks have some non-overlapping samples, systems usually perform similarly on them.
@@
 Otter $F\!\!\to\!\!P$ e-Otter $F\!\!\to\!\!P$ Change Otter $F\!\!\to\!\!P$ e-Otter $F\!\!\to\!\!P$ Change
-TDD-Bench
-Verified (449
-samples)
-Planner 145 (32.3%) 179 (39.9%) 23.4% 141 (31.4%) 167 (37.2%) 18.4%
-Context
-Masking
-full 125 (27.8%) 147 (32.7%) 17.6% 110 (24.5%) 135 (30.1%) 22.7%
+TDD-Bench Verified (449 samples) Planner 145 (32.3%) 179 (39.9%) 23.4% 141 (31.4%) 167 (37.2%) 18.4%
+Context Masking full 125 (27.8%) 147 (32.7%) 17.6% 110 (24.5%) 135 (30.1%) 22.7%
 testLoc 117 (26.1%) 130 (29.0%) 11.1% 115 (25.6%) 135 (30.1%) 17.4%
@@
 none 100 (22.3%) 127 (28.3%) 27.0% 110 (24.5%) 137 (30.5%) 24.5%
-Issue
-Morphing
-standard 137 (30.5%) 183 (40.8%) 33.6% 104 (23.2%) 135 (30.1%) 29.8%
+Issue Morphing standard 137 (30.5%) 183 (40.8%) 33.6% 104 (23.2%) 135 (30.1%) 29.8%
 simple 137 (30.5%) 182 (40.5%) 32.8% 104 (23.2%) 128 (28.5%) 23.1%
@@
 initPatch 151 (33.6%) 191 (42.5%) 26.5% 119 (26.5%) 149 (33.2%) 25.2%
-SWT-bench
-Lite (276
-samples)
-Planner 83 (30.1%) 99 (35.9%) 19.3% 70 (25.4%) 80 (29.0%) 14.3%
-Context
-Masking
-full 67 (24.3%) 89 (32.2%) 32.8% 58 (21.0%) 71 (25.7%) 22.4%
+SWT-bench Lite (276 samples) Planner 83 (30.1%) 99 (35.9%) 19.3% 70 (25.4%) 80 (29.0%) 14.3%
+Context Masking full 67 (24.3%) 89 (32.2%) 32.8% 58 (21.0%) 71 (25.7%) 22.4%
 testLoc 67 (24.3%) 82 (29.7%) 22.4% 55 (19.9%) 76 (27.5%) 38.2%
@@
 none 57 (20.7%) 68 (24.6%) 19.3% 58 (21.0%) 75 (27.2%) 29.3%
-Issue
-Morphing
-standard 80 (29.0%) 95 (34.4%) 18.8% 62 (22.5%) 76 (27.5%) 22.6%
+Issue Morphing standard 80 (29.0%) 95 (34.4%) 18.8% 62 (22.5%) 76 (27.5%) 22.6%
 simple 68 (24.6%) 88 (31.9%) 29.4% 57 (20.7%) 69 (25.0%) 21.1%
@@
 How good are LLMs as critics?
-We just saw that that critics play an important role in test repair.
+We just saw that critics play an important role in test repair.
 But how good are LLMs as critics?
@@
 Morphs and masks yield heterogeneous prompts, which in turn increase the diversity of candidate tests, thus hopefully increasing $F\!\!\to\!\!P$ @ N.
-We reuse 5 heterogeneous prompts including masks from Otter++ (ahmed_et_al_2025) and add 5 more prompts using morphs for a total of 10 heterogeneous prompts.
+We reuse 5 heterogeneous prompts including masks from Otter++ (Ahmed et al., 2025) and add 5 more prompts using morphs for a total of 10 heterogeneous prompts.
 Table 6 shows that morphs increase $F\!\!\to\!\!P$ @ N, even without masks, both before (Otter) and after (e-Otter) test repair.
@@
 Benchmark Model Category N Otter $F\!\!\to\!\!P$ @ N Change e-Otter $F\!\!\to\!\!P$ @ N Change
-TDD-Bench
-Verified
-Claude-3.7-
-Sonnet
-plan + mask 5 210 (46.8%) - 260 (57.9%) -
+TDD-Bench Verified Claude-3.7- Sonnet plan + mask 5 210 (46.8%) - 260 (57.9%) -
 plan + morph 6 255 (56.8%) 21.4% 305 (67.9%) 17.3%
@@
 plan + mask + morph 10 237 (52.8%) 20.3% 260 (57.9%) 18.7%
-SWT-bench
-Lite
-Claude-3.7-
-Sonnet
-plan + mask 5 113 (40.9%) - 146 (52.9%) -
+SWT-bench Lite Claude-3.7- Sonnet plan + mask 5 113 (40.9%) - 146 (52.9%) -
 plan + morph 6 141 (51.1%) 24.8% 162 (58.7%) 11.0%
@@
 Why not use high temperature instead of heterogeneous prompting?
-Prior works (xia_et_al_2025; ehrlich_et_al_2025; chen2021evaluating) mostly depend on high temperature for multi-sampling.
+Prior works (Xia et al., 2025; Ehrlich et al., 2025; Chen et al., 2021) mostly depend on high temperature for multi-sampling.
 Figure 5 compares the results of heterogeneous prompting with the samples generated with high temperature.
-We tried three distinct temperatures (0.5, 0.8, and 1.0) and find that the best results are obtained at a temperature of 0.8, which is consistent with prior research (chen2021evaluating).
+We tried three distinct temperatures (0.5, 0.8, and 1.0) and find that the best results are obtained at a temperature of 0.8, which is consistent with prior research (Chen et al., 2021).
 However, our approach, “plan + mask + morph”, significantly outperforms samples generated with high temperature.
@@
 Each prompt contributes between 0.3% and 4.6% to $F\!\!\to\!\!P$ @ N.
-Even though some prompts do not add many new instances, increasing $F\!\!\to\!\!P$ @ N is still beneficial, as if gives the test selector a larger pool of valid candidates to choose from, increasing the chances of selecting the most suitable ones.
+Even though some prompts do not add many new instances, increasing $F\!\!\to\!\!P$ @ N is still beneficial, as it gives the test selector a larger pool of valid candidates to choose from, increasing the chances of selecting the most suitable ones.
 Although we report most results using two closed-source models, Claude-3.7-Sonnet and GPT-4o, we also investigate the effectiveness of execution-augmented test repair and heterogeneous prompting on a relatively weaker open-source model, Mistral-large (123 billion parameters).
@@
 This indicates that our approaches are also applicable to open-source models.
-Table 8 shows how much a morph can increase $F\!\!\to\!\!P$ @ 2 from $F\!\!\to\!\!P$ @ 1 (using base e-Otter) for TDD-bench-Verified because multiple morphs can independently benefit a single issue.
-Each morph increases $F\!\!\to\!\!P$ by 29%-32% for Claude and 11%-21% for GPT-4o.
-Table 8. Contribution of each prompt to $F\!\!\to\!\!P$ @ N
-Variants Claude-3.7-Sonnet GPT-4o
-$F\!\!\to\!\!P$ @ 2 Change $F\!\!\to\!\!P$ @ 2 Change
-Standard 235 +31.28% 194 +16.16%
-Simple 235 +31.28% 186 +11.37%
-dropCode 231 +29.05% 199 +19.16%
-initTest 228 +27.37% 197 +17.96%
-initPatch 236 +31.84% 203 +21.55%
+Multiple morphs can independently benefit a single issue on TDD-bench-Verified.
+Using the base e-Otter setup, applying an additional morph increases $F\!\!\to\!\!P$ @ 2 relative to $F\!\!\to\!\!P$ @ 1 by 29%–32% for Claude and 11%–21% for GPT-4o.
 Finding 3. The samples with “plan + mask + morph” increase $F\!\!\to\!\!P$ @ N by more than 13% compared to “plan + mask”, regardless of the model and benchmark.
@@
 Benchmark Model Pre-filter Selection $F\!\!\to\!\!P$
-TDD-Bench
-Verified
-Clade-3.7-Sonnet None Random 200.5 (44.7%)
+TDD-Bench Verified Clade-3.7-Sonnet None Random 200.5 (44.7%)
 None Coverage 213.0 (47.4%)
@@
 CodePatch Coverage 231.0 (51.4%)
-SWT-bench
-Verified
-Clade-3.7-Sonnet None Random 100.9 (36.5%)
+SWT-bench Verified Clade-3.7-Sonnet None Random 100.9 (36.5%)
 None Coverage 111.0 (40.2%)
@@
 Ahmed et al.
-(ahmed_et_al_2025) took the top 22 systems from the leaderboard and ran the five tests generated by Otter++.
+(Ahmed et al., 2025) took the top 22 systems from the leaderboard and ran the five tests generated by Otter++.
 They filtered out a code patch if all the tests failed on it.
@@
 Mündler et al.
-(mundler_et_al_2024) performed a similar experiment with one system.
+(Mündler et al., 2024) performed a similar experiment with one system.
 Note that neither of these two prior approaches could achieve a $F\!\!\to\!\!P$ rate beyond 50% even with 5 samples.
@@
 To show that our proposed approach is effective for non-contaminated models, we extend our evaluation on SWE-rebench with the Claude-3.7-Sonnet model.
-The training data cut-off date for Claude-3.7-Sonnet is November 2024 (claude-cutoff).
-To address data contamination, we consider instances from SWE-rebench (badertdinov2025swe) with a creation date in 2025 and check whether the developers’ written golden tests go from fail-to-pass to select the instances.
+The training data cut-off date for Claude-3.7-Sonnet is November 2024 (3).
+To address data contamination, we consider instances from SWE-rebench (Badertdinov et al., 2025) with a creation date in 2025 and check whether the developer-written golden tests go from fail-to-pass to select the instances.
 This experiment uses 267 instances.
@@
 The absolute numbers on the non-contaminated SWE-rebench are notably lower than on TDD-Bench or SWT-bench, and contamination might be a major factor behind that.
-Another factor is that the samples from SWE-rebench come from 152 projects, most of which are less popular as the projects in SWE-bench.
-Popular projects are relatively older, larger, and more likely to be included in training data.
+Another factor is that the samples from SWE-rebench come from 152 projects, most of which are less popular than the projects in SWE-bench.
+Popular projects are relatively older, larger, and may be more likely to have well-written issue descriptions.
 Despite lower absolute numbers, our approach shows a great relative improvement, demonstrating that our two primary contributions also work for contamination-free benchmarks like SWE-rebench.
@@
 5. Discussion
-This section briefly discusses model contamination, test coverage, and a detailed analysis of the tests.
+This section discusses model contamination, test coverage, and a detailed analysis of the tests.
 Contamination through Issue Morphs.
@@
 Both prior works by Ahmed et al.
-(ahmed_et_al_2025) and Mündler et al.
-(mundler_et_al_2024) addressed this issue and show that the model-generated tests are very different from developer-written tests and the impact of contamination is minimal.
+(Ahmed et al., 2025) and Mündler et al.
+(Mündler et al., 2024) addressed this issue and show that the model-generated tests are very different from developer-written tests and the impact of contamination is minimal.
 However, e-Otter++ might potentially introduce new contamination.
@@
 Since we have four raters, it is infeasible to do majority voting on the rated instances.
-Instead, we adapted a stricter metric to assign the final label.
+Instead, we adopted a stricter metric to assign the final label.
 If at least two raters think there is some addition, removal, or contamination, we counted it as “1”, otherwise “0”.
 Table 13 shows the inter-rater agreement and outcome of the study.
-We use Fleiss’ Kappa (fleiss1971measuring) to compute inter-rater agreement, and the raters show fair to substantial agreement [^3] Kappa ¡ 0: Poor, 0.01–0.20: Slight, 0.21–0.40: Fair, 0.41–0.60: Moderate, 0.61–0.80: Substantial, 0.81–1.00: Almost perfect agreement..
+We use Fleiss’ Kappa (Fleiss, 1971) to compute inter-rater agreement, and the raters show fair to substantial agreement [^3] Kappa ¡ 0: Poor, 0.01–0.20: Slight, 0.21–0.40: Fair, 0.41–0.60: Moderate, 0.61–0.80: Substantial, 0.81–1.00: Almost perfect agreement..
 Note that for contamination, we observe a very low kappa value because all raters mostly choose the same category.
-This phenomenon is known as the kappa paradox (feinstein1990high).
+This phenomenon is known as the kappa paradox (Feinstein and Cicchetti, 1990).
 Overall, 40% of modified issues have some information addition and 68% of issues have some information removal.
@@
 Only for 1 (2%) sample out of 50 samples, more than one rater think the rewriting might have caused some contamination.
-Therefore, we can infer that issue morph does not contaminate the issue.
-We share the ratings from each individual rater in the supplementary material.
+Therefore, we can infer that issue morphs do not contaminate the issue.
 Table 13. Manual validation to investigate model contamination or memorization
@@
 Both 50 0.60 Moderate 34 (68%)
-Contamination Claude-3.7-Sonnet 25
-Kappa value is not meaningful
-because all raters mostly choose
-the same category.
-0 (0%)
+Contamination Claude-3.7-Sonnet 25 Kappa value is not meaningful because all raters mostly choose the same category. 0 (0%)
 GPT-4o 25 1 (4%)
@@
 Figure 7 shows that $F\!\!\to\!\!P$ tests have higher coverage compared to other tests.
-These findings are consistent with prior works (ahmed_et_al_2025; mundler_et_al_2024) and motivate using coverage as a tie-breaker for test selection (of course, test selection uses coverage on generated code patches, not golden ones).
+These findings are consistent with prior works (Ahmed et al., 2025; Mündler et al., 2024) and motivate using coverage as a tie-breaker for test selection (of course, test selection uses coverage on generated code patches, not golden ones).
 Figure 7. Coverage of the generated tests
@@
 As a result, the findings may not generalize to other programming languages.
-However, it is worth noting that prior works (jimenez_et_al_2024; xia_et_al_2025; zhang_et_al_2024), which share this same limitation, have still been impactful and contributed significantly to the field.
-One of the key concerns with SWE-bench related work is the contamination or memorization problem of the model.
-To address this problem, we evaluated e-Otter++ on SWE-rebench and observed similar impact from execution-augmented test repair and heterogeneous prompting.
-Another limitation of e-Otter++ is that it considers only one test file and generates only one block of code.
+However, prior works (Jimenez et al., 2024; Xia et al., 2025; Zhang et al., 2024) with the same limitation have still been impactful and contributed significantly to the field.
+One of the key concerns with SWE-bench related work is the contamination or memorization problem of the model (Badertdinov et al., 2025; Liang et al., 2025a).
+To address this problem, we evaluated e-Otter++ on SWE-rebench (Badertdinov et al., 2025) and observed similar impact from execution-augmented test repair and heterogeneous prompting.
+(We did not use REPOCOD (Liang et al., 2025b), since it is not contamination-free, nor LiveCodeBench (Jain et al., 2025a), since it is not repository-level.) Another limitation of e-Otter++ is that it considers only one test file and generates only one block of code.
 In real-world projects, test code can be spread across multiple files or blocks of code.
@@
 These benchmarks allow us to apply the golden patch (written by real developers and likely to be correct), run the test, and check both how well it covers the edited lines and whether it goes from fail to pass.
-While this is no correctness guarantee, it is a strong proxy for correctness and followed by prior works (mundler_et_al_2024; ahmed_et_al_2025).
-Besides, using imperfect patches or tests comes with certain risks, but prior work (chen_et_al_2023) has shown that dual execution agreement can help discard samples with potential errors.
-To further mitigate risk, we can use separate systems for surrogate patches and the proposed final patch (of course, with increased cost).
+While this is no correctness guarantee, it is a strong proxy for correctness and followed by prior works (Mündler et al., 2024; Ahmed et al., 2025).
+Besides, using imperfect patches or tests comes with certain risks, but prior work (Chen et al., 2023) has shown that dual execution agreement can help discard samples with potential errors.
+To further mitigate risk, one could use separate systems for surrogate patches and the proposed final patch (of course, with increased cost).
 7. Related Work
 There has been a flurry of recent papers about generating reproduction tests from issues, which is the same problem we tackle in our paper.
-Libro generates tests for Java bug reports (kang_yoon_yoo_2023).
+Libro generates tests for Java bug reports (Kang et al., 2023).
 It repairs tests based on compiler feedback, not execution feedback like we do, and generates several tests using high LLM decoding temperature, not morphs and masks like we do.
 Plein et al.
-also explore the feasibility of generating tests for Java bug reports, albeit without test repair (plein_et_al_2024).
+also explore the feasibility of generating tests for Java bug reports, albeit without test repair (Plein et al., 2024).
 Mündler et al.
-(mundler_et_al_2024) introduce SWE-Agent+, a Python reproduction test generation agent derived from the SWE-Agent (yang_et_al_2024) code patch generator.
+(Mündler et al., 2024) introduce SWE-Agent+, a Python reproduction test generation agent derived from the SWE-Agent (Yang et al., 2024) code patch generator.
 While SWE-Agent+ has access to tools that can give it execution feedback, it leverages such feedback less effectively than e-Otter++.
-The same holds for EvoCoder (lin_et_al_2024), also derived from SWE-Agent, and for USEAgent (applis_et_al_2025), derived from AutoCodeRover (zhang_et_al_2024).
-AEGIS is a novel multi-agent system for reproduction test generation (wang_et_al_2024).
+The same holds for EvoCoder (Lin et al., 2024), also derived from SWE-Agent, and for USEAgent (Applis et al., 2025), derived from AutoCodeRover (Zhang et al., 2024).
+AEGIS is a novel multi-agent system for reproduction test generation (Wang et al., 2024).
 It repairs tests based on execution feedback from $c_{\textrm{old}}$, but unlike our work, does not augment the test repair prompt.
 Also, it standardizes issue descriptions, but unlike our work, does not use morphs or masks to increase test candidate pool diversity.
-Otter++ (ahmed_et_al_2025) is an LLM-based workflow for reproduction test generation that we use as a component in e-Otter++, which extends it in three important ways: (i) augmenting the test repair prompt; (ii) using morphs (Otter++ only uses masks); and (iii) leveraging execution feedback from generated candidate code patches.
+Otter++ (Ahmed et al., 2025) is an LLM-based workflow for reproduction test generation that we use as a component in e-Otter++, which extends it in three important ways: (i) augmenting the test repair prompt; (ii) using morphs (Otter++ only uses masks); and (iii) leveraging execution feedback from generated candidate code patches.
 Our experiments show that these novel contributions greatly improve the $F\!\!\to\!\!P$ rate.
-BRT Agent is a new agent for generating reproduction tests (cheng_et_al_2025), and Issue2Test is a new non-agentic workflow for the same task (nashid_et_al_2025); neither augments test repair prompts, uses masks or morphs, or leverages candidate code patches.
+BRT Agent is a new agent for generating reproduction tests (Cheng et al., 2025), and Issue2Test is a new non-agentic workflow for the same task (Nashid et al., 2025); neither augments test repair prompts, uses masks or morphs, or leverages candidate code patches.
 Unlike e-Otter++ and the above papers on using generated code patches to help pick or improve reproduction tests, there is also work on the opposite: using reproduction tests to help pick or improve code patches.
-Papers on using tests to help pick (but not improve) code patches include Smith et al.’s 2015 paper (smith_et_al_2015), CodeT (chen_et_al_2023), Agentless (xia_et_al_2025), and R2E-Gym (jain_et_al_2025).
-Similarly, several papers use tests to help improve (but not pick) code patches (zhang_et_al_2024; yang_et_al_2024; chen_et_al_2024; arora_et_al_2024; ruan_zhang_roychoudhury_2024; mathews_nagappan_2024).
-Some papers even do both, using tests to help pick and improve code patches: CodeMonkeys (ehrlich_et_al_2025) runs several parallel workflows, each of which iteratively refines a $\langle c_{\mathrm{new}},y\rangle$ pair; and PatchPilot (li_et_al_2025) uses reproduction tests for bug localization before code patch generation, and later uses them again to pick among multiple candidate code patches.
+Papers on using tests to help pick (but not improve) code patches include Smith et al.’s 2015 paper (Smith et al., 2015), CodeT (Chen et al., 2023), Agentless (Xia et al., 2025), and R2E-Gym (Jain et al., 2025b).
+Similarly, several papers use tests to help improve (but not pick) code patches (Zhang et al., 2024; Yang et al., 2024; Chen et al., 2024; Arora et al., 2024; Ruan et al., 2024; Mathews and Nagappan, 2024).
+Some papers even do both, using tests to help pick and improve code patches: CodeMonkeys (Ehrlich et al., 2025) runs several parallel workflows, each of which iteratively refines a $\langle c_{\mathrm{new}},y\rangle$ pair; and PatchPilot (Li et al., 2025) uses reproduction tests for bug localization before code patch generation, and later uses them again to pick among multiple candidate code patches.
 However, while most of the work described in this paragraph uses generated tests, test generation is not their main focus: their reproduction test generation components are more basic, and they lack empirical evaluation of the $F\!\!\to\!\!P$ rate of generated tests.
 That said, they demonstrate a widespread need for our work.
+Besides, Chen et al.
+(Chen et al., 2025a) show that, together with regression tests, bug reproduction tests can improve the performance of issue resolution systems.
 8. Conclusion
@@
 Overall, we hope that by automatically generating reproduction tests, we can take the tedium out of test-driven development (TDD), so more developers and software projects can reap the benefits of TDD.
-All our generated tests and logs are available at https://doi.org/10.5281/zenodo.15860535.
+All our generated tests and logs are available in our repository (2).
 Acknowledgements.
-We gratefully acknowledge Ibragim Badertdinov and Chunqiu Steven Xia for sharing their model-generated SWE-rebench and Agentless patches.
+We gratefully acknowledge Ibragim Badertdinov and Chunqiu Steven Xia for sharing SWE-rebench and Agentless patches.